### v1

# Another look at inference after prediction

Abstract
Prediction-based (PB) inference is increasingly used in applications where the outcome of interest is difficult to obtain, but its predictors are readily available. Unlike traditional inference, PB inference performs statistical inference using a partially observed outcome and a set of covariates by leveraging a prediction of the outcome generated from a machine learning (ML) model. Motwani and Witten (2023) recently revisited two innovative PB inference approaches for ordinary least squares. They found that the method proposed by Wang et al. (2020) yields a consistent estimator for the association of interest when the ML model perfectly captures the underlying regression function. Conversely, the prediction-powered inference (PPI) method proposed by Angelopoulos et al. (2023a) yields valid inference regardless of the model’s accuracy. In this paper, we study the statistical efficiency of the PPI estimator. Our analysis reveals that a more efficient estimator, proposed 25 years ago by Chen and Chen (2000), can be obtained by simply adding a weight to the PPI estimator. We also contextualize PB inference with methods from the economics and statistics literature dating back to the 1960s. Our extensive theoretical and numerical analyses indicate that the Chen and Chen (CC) estimator offers a balance between robustness to ML model specification and statistical efficiency, making it the preferred choice for use in practice.
Keywords: Missing data, prediction-based inference, semi-parametric inference
1 Introduction
In the rapidly evolving landscape of data science, accurately predicting an outcome is often just an initial step in scientific inquiry. Researchers are increasingly turning to machine learning (ML) models to obtain predictions for outcomes that are prohibitively time-consuming or expensive to directly measure, and then drawing statistical inferences from the ML model’s predictions (Hoffman et al., 2024; Fan et al., 2024; Gao et al., 2024; Miao and Lu, 2024). This prediction-based (PB) inference scenario was introduced by Wang et al in the context of global health and genomics applications, and is comprised of two steps outlined in Figure 1 (Wang et al., 2020). First, the outcome of interest is imputed using a prediction from a pre-trained model. Then, the predictions are used to draw statistical inference on the parameter of interest, such as the association between an outcome and a set of covariates.
Figure 1: Prediction-based (PB) inference. $Y$ is the outcome of interest, which is only partially observed, ${\bf{Z}}$ is a set of predictors of $Y$, $\widehat{Y}$ is the prediction of $Y$ based on ${\bf{Z}}$, and ${\bf{X}}$ is a set of covariates. The focus of scientific inquiry is on the association between ${\bf{X}}$ and $Y$. Traditional statistical inference relies solely on the data on ${\bf{X}}$ and $Y$ while PB inference incorporates $\widehat{Y}$ to maximize the use of all available data. Dark red represents labeled observations while light red represents unlabeled observations.
In the last several years, applications of PB inference have arisen in diverse fields, spanning genetics, biology, and medicine. For example, population biobanks, such as the UK Biobank ($N\approx$ 500K individuals), provide an unprecedented data source for high-throughput genetic research. However, many traits and disease outcomes are only partially observed due to the time and cost of measurement. ML models are commonly used to predict missing outcomes based on fully observed baseline data and the predictions are utilized for genetic discovery (Tong et al., 2020; Alipanahi et al., 2021; Cosentino et al., 2023; Dahl et al., 2023; An et al., 2023; McCaw et al., 2024; Liu et al., 2024; Miao et al., 2024). Similarly, in the field of structural biology, AlphaFold provides highly accurate structural predictions without the need for expensive and labor-intensive experiments for proteomic studies (Jumper et al., 2021). In the analysis of electronic health records data, large language models can rapidly ascertain disease status at an accuracy competitive with manual chart review by experts, enabling downstream epidemiological and clinical studies (Yang et al., 2022; Zhou et al., 2022; Alsentzer et al., 2023).
Given the relevance of PB inference, Motwani and Witten (2023) recently revisited two innovative methods in this area. In the context of least squares estimation, they showed that the approach proposed by Wang et al. (2020) provides a consistent estimator for the association between an outcome and a set of covariates only in restrictive scenarios. For example, when the pre-trained model perfectly captures the regression function. This is a significant practical limitation as it is essential that PB inference methods yield valid inference in the likely scenario that the prediction model is imperfect. In contrast, the prediction powered inference (PPI) approach proposed by Angelopoulos et al. (2023a) possesses this necessary property and yields valid inference regardless of the quality of the model.
Here we build on the work on Motwani and Witten to study the statistical efficiency of the PPI method. We highlight that a more efficient estimator can be obtained by simply adding a weight to the PPI estimator, an approach that was proposed by Chen & Chen (CC) more than 25 years ago in the context of double sampling designs (Chen and Chen, 2000). We begin by formally introducing the PB inference problem and reviewing the contributions of Motwani and Witten in Section 2. We then study the asymptotic properties of the PPI estimator and introduce the CC estimator in Section 3. Our key finding is that the CC estimator is optimal with respect to statistical efficiency in a class of estimators to which the PPI estimator belongs. We also point to earlier literature from the statistics and economics literature dating back to the 1960s that provides important historical context for the PB inference problem in Section 4. In Section 5, we evaluate our findings with extensive simulation studies. Our results show that the CC estimator offers a balance between robustness to ML model specification and statistical efficiency, making it the preferred choice for use in practice. We close with additional discussion in Section 6.
2 Problem setting & background
2.1 Notation and Assumptions
We let $Y$ denote the outcome of interest and ${\bf{Z}}$ a set of predictors. $Y$ is only partially observed and a pre-trained ML model is available to predict it on the basis of ${\bf{Z}}$. The ML model, $\hat{f}:\mathcal{Z}\rightarrow\mathcal{Y}$, maps from the predictor space, $\mathcal{Z}$, to the outcome space, $\mathcal{Y}$. The predictions of $Y$ are denoted as $\widehat{Y}=\hat{f}({\bf{Z}})$. We let ${\bf{X}}=(1,X_{1},\dots,X_{p})^{{\sf\scriptscriptstyle{T}}}$ be a set of covariates that are potentially associated with $Y$ and that are the focus of scientific inquiry. ${\bf{X}}$ may also be predictive of $Y$ and these covariates may or may not be included in ${\bf{Z}}$. We assume that ${\bf{X}}$ and ${\bf{Z}}$ are both fully observed. The available data therefore consists of a labeled dataset and an independent unlabeled dataset, respectively denoted as
=$\mathscr{L}=\{(Y_{i},{\bf{Z}}_{i}^{{\sf\scriptscriptstyle{T}}},{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}})^{{\sf\scriptscriptstyle{T}}}\mid i=1,\dots,n_{\text{\tiny{lab}}}\}\quad\mbox{and}\quad\mathscr{U}=\{({\bf{Z}}_{i}^{{\sf\scriptscriptstyle{T}}},{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}})^{{\sf\scriptscriptstyle{T}}}\mid i=n_{\text{\tiny{lab}}}+1,\dots,n\},$
where $n=n_{\text{\tiny{lab}}}+n_{\text{\tiny{unlab}}}$, $n_{\text{\tiny{unlab}}}$ is the number of unlabeled observations, and $n_{\text{\tiny{lab}}}/n\to\pi$ as $n\to\infty$ for $\pi\in(0,1)$.
We focus on the case in which $\mathscr{L}$ and $\mathscr{U}$ are independent and identically distributed samples from the population of interest. More complex scenarios have been the focus of recent work (Angelopoulos et al., 2023a; Zrnic and Candès, 2024a). Additionally, $\hat{f}$ is assumed to be pre-trained with a dataset that is independent of $\mathscr{L}$ and $\mathscr{U}$ and is therefore treated as “fixed” (i.e., the variation in estimating $\hat{f}$ is not of interest in downstream inference). In fact, $\hat{f}$ may be a black box function in that its training data and specification are not be available to the users. Approaches that address uncertainty in the estimation of $\hat{f}$ have recently been proposed in Zrnic and Candès (2024b).
2.2 Parameter of interest
Following Motwani and Witten (2023), we consider least squares estimation and aim to make inference on the population parameter
$\bm{\beta}^{*}\coloneq\operatorname*{arg\,min}_{\bm{\beta}}\mathbb{E}[(Y-{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\beta})^{2}]=[\mathbb{E}({\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}})]^{-1}\mathbb{E}({\bf{X}}Y),$
which quantifies the association between $Y$ and ${\bf{X}}$. Throughout, we use superscripts to denote different estimators of $\bm{\beta}^{*}$. For example, the simplest unbiased estimator of $\bm{\beta}^{*}$, denoted as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, uses only $\mathscr{L}$ for estimation and is the solution to the estimating equation
$\sum_{i=1}^{n_{\text{\tiny{lab}}}}\bm{\phi}(Y_{i},{\bf{X}}_{i};\bm{\beta})\coloneq\sum_{i=1}^{n_{\text{\tiny{lab}}}}{\bf{X}}_{i}(Y_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\bm{\beta})=\bm{0}.$
2.3 Motivation for PB inference
The goal of PB inference is to utilize the abundant realizations of $\widehat{Y}$, which can be computed in the unlabeled data, to provide valid and efficient inference on the association between $Y$ and ${\bf{X}}$. Motwani and Witten consider the question of validity and we study efficiency in Section 3. With respect to validity, PB inference should enable users to draw correct conclusions regarding the magnitude and direction of $\bm{\beta}^{*}$ and to perform reliable uncertainty quantification. Incorporating information from $\widehat{Y}$ should not change the parameter being estimated and the credibility of inferences should not depend on the quality of predictions from $\hat{f}$.
The need for PB inference is highlighted by limitations of “naive” inference, in which inference on $\widehat{Y}$ is performed in lieu of inference on $Y$ (Wang et al., 2020). The parameter targeted by naive inference is
$\bm{\gamma}^{*}\coloneq\operatorname*{arg\,min}_{\bm{\gamma}}\mathbb{E}[(\widehat{Y}-{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma})^{2}]=[\mathbb{E}({\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}})]^{-1}\mathbb{E}({\bf{X}}\widehat{Y}).$
A least squares estimator for $\bm{\gamma}^{*}$, denoted as $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$, is the solution to
$\sum_{i=n_{\text{\tiny{lab}}}+1}^{n}\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma})\coloneq\sum_{i=n_{\text{\tiny{lab}}}+1}^{n}{\bf{X}}_{i}(\widehat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma})=\bm{0}.$
Critically, $\bm{\gamma}^{*}$ bears no necessary correspondence with $\bm{\beta}^{*}$ except under highly restrictive scenarios, such as when ${\bf{Z}}$ contains ${\bf{X}}$ and the ML model captures the true regression function. The fundamental problem is that $(\widehat{Y},{\bf{X}}^{{\sf\scriptscriptstyle{T}}})^{{\sf\scriptscriptstyle{T}}}$ has a different distribution than $(Y,{\bf{X}}^{{\sf\scriptscriptstyle{T}}})^{{\sf\scriptscriptstyle{T}}}$. To overcome these limitations, PB inference methods utilize $\mathscr{L}$ together with $\mathscr{U}$ to essentially balance the relative contributions of $Y$ and $\widehat{Y}$ in making inference on $\bm{\beta}^{*}$.
2.4 PB inference methods
Motwani and Witten revisited two innovative strategies for PB inference. The first is the proposal of Wang et al. (2020), which is a two-step procedure to incorporate $\widehat{Y}$ into estimation. First, $\mathscr{L}$ is used to model the relationship between $Y$ and $\widehat{Y}$ to address potential inaccuracies in the predictions. Then, this model is used together with $\mathscr{U}$ to make inference on $\bm{\beta}^{*}$, with analytical or bootstrap approaches described. In their illuminating work, Motwani and Witten show that the estimator of Wang et al is consistent for $\bm{\beta}^{*}$ under the strong and unrealistic assumption that the ML model captures the true regression function. In other scenarios, the Wang et al estimator may not be consistent for $\bm{\beta}^{*}$ even while its standard error converges to 0 as $(n_{\text{\tiny{lab}}},n_{\text{\tiny{unlab}}})\to\infty$. Consequently, the familiar Wald statistic will not converge in distribution under the null, resulting in type I error that is not controlled at the nominal level and confidence intervals with inadequate coverage.
Motwani and Witten also study the PPI method, which has the crucial property that inference remains valid regardless of the quality of the predictions. Stated otherwise, the PPI method does not require that $\hat{f}$ capture the true regression function for the estimator to maintain consistency or to yield reliable hypothesis testing and interval estimation. The proposed estimator utilizes a common statistical technique involving the augmentation of an initial consistent estimator for the parameter of interest with an additional term that converges in probability to 0. This approach dates back to at least Robins et al. (1994), and has been applied in multiple settings to improve estimation efficiency outside of the PB inference problem (e.g., Chen and Chen (2000); Tsiatis (2006); Tian et al. (2012); Wang and Wang (2015)). We return to this point in Section 4.
The PPI estimator is
$\hat{\bm{\beta}}^{\text{\tiny{PPI}}}=\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}\right),$
where $\hat{\bm{\gamma}}^{\text{\tiny{lab}}}$ is the solution to $\sum_{i=1}^{n_{\text{\tiny{lab}}}}\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma})=\boldsymbol{0}$ [^1] Angelopoulos et al. (2023a) express $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ as $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}-\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\beta}}^{\text{\tiny{lab}}}\right)$ where $\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\beta}}^{\text{\tiny{lab}}}\right)$ is referred to as a “rectifier.” The rectifier serves the purpose of correcting potential bias in $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ due to inaccuracies in $\widehat{Y}$ for $Y$. We reframe the PPI method to study its statistical efficiency.. In contrast to the approach of Wang et al, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is consistent for $\bm{\beta}^{*}$ and $n^{\frac{1}{2}}(\hat{\bm{\beta}}^{\text{\tiny{PPI}}}-\bm{\beta}^{*})$ converges weakly to a zero-mean multivariate normal distribution regardless of the quality of the prediction. The proposed method therefore asymptotically controls the type I error for Wald-based hypothesis testing and provides confidence intervals with the nominal coverage level.
3 Efficiency analysis
As PB inference methods suppose the availability of at least a small labeled data set, a natural baseline estimator for comparison is $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, which simply regresses $Y$ on ${\bf{X}}$ within $\mathscr{L}$. For a PB inference method to be preferred over this baseline, it should not only yield valid inference, but also improve statistical efficiency relative to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when the prediction is informative of $Y$, and perform no worse than estimation based on $\mathscr{L}$ when $\widehat{Y}$ is uninformative of $Y$. For the remainder of this paper, we study the latter property and analyze an estimator very similar to $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ given by
$\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}=\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}\right),$
where $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$ is the solution to $\sum_{i=1}^{n}\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma})=\boldsymbol{0}$. The only difference between $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is that all of the data, rather than the unlabeled data, is used to estimate $\bm{\gamma}^{*}$ in the augmentation term (i.e., $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ is replaced by $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$). When $\pi\approx 0$, $\hat{\bm{\gamma}}^{\text{\tiny{all}}}\approx\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ so that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}\approx\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$. In general, it can be shown that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ gains efficiency over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ whenever $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ gains over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ can be more efficient than $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ when $\widehat{Y}$ is informative of $Y$ (see Appendix B, specifically Section B.3, for a detailed analysis). Additionally, estimators of this form are commonly used in the semi-parametric inference literature, which we draw upon in the subsequent section in our efficiency analysis.
3.1 Asymptotic analysis of the PPI estimator
To study the asymptotic variance of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, Theorem 1 provides the influence function expansion of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$. All of the asymptotic results for $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ are detailed in Appendix A.
Theorem 1 (Influence function expansion of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$)
Let $R_{i}$ be the indicator of whether the $i$th observation is labeled. Then,
$n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}-\bm{\beta}^{*}\right)=n^{-\frac{1}{2}}\sum_{i=1}^{n}{\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})+o_{p}(1),$
where ${\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})={\bf{A}}\left\{\left[\ \bm{\phi}(Y_{i},{\bf{X}}_{i};\bm{\beta}^{*})\right]\frac{R_{i}}{\pi}+\left[\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma}^{*})\right]\frac{\pi-R_{i}}{\pi}\right\}$ and ${\bf{A}}=[\mathbb{E}({\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}})]^{-1}$. Therefore, $n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}-\bm{\beta}^{*}\right)$ converges weakly to a zero-mean multivariate normal distribution with variance-covariance
$\mathbb{E}\left[{\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})\hskip 2.84526pt{\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})^{{\sf\scriptscriptstyle{T}}}\right].$
We use Theorem 1 to compare the asymptotic variance-covariance of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ with that of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. Corollary 1.1 shows that the benefit of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ scales directly with the predictiveness of $\widehat{Y}$ for $Y$ and inversely with the proportion of labeled data, $\pi$.
Corollary 1.1 (Comparison of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ with $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$)
Let $\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}=\mathbb{E}\left(\epsilon_{\widehat{Y}}^{2}{\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\right)$, $\boldsymbol{\Sigma}_{Y,\widehat{Y}}=\mathbb{E}\left(\epsilon_{Y}\epsilon_{\widehat{Y}}{\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\right)$, $\epsilon_{Y}=Y-{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\beta}^{*}$ and $\epsilon_{\widehat{Y}}=\widehat{Y}-{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma}^{*}$. The difference in the asymptotic variance-covariance (aVar) of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is
$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})$ $=n\left[\text{aVar}(\hat{\bm{\beta}}^{\text{\tiny{lab}}})-\text{aVar}(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})\right]=\left(\frac{1}{\pi}-1\right){\bf{A}}\left(2\boldsymbol{\Sigma}_{Y,\widehat{Y}}-\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}\right){\bf{A}}.$
$\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is therefore more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ if and only if $2\boldsymbol{\Sigma}_{Y,\widehat{Y}}\succ\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}$. A sufficient condition for $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ to have lower variance than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ is that
$\mathbb{E}\left[\left(\epsilon_{\widehat{Y}}-\epsilon_{Y}\right)^{2}\mid{\bf{X}}\right]<\mathbb{E}\left(\epsilon_{Y}^{2}\mid{\bf{X}}\right).$
More specifically, the sufficient condition in Corollary 1.1 coincides with the intuition that incorporating unlabeled data into estimation can improve efficiency when $\widehat{Y}$ has additional predictive power for $Y$ beyond what is already explained by the regression model of $Y$ on ${\bf{X}}$. This condition is similar to the result for mean estimation provided in the Supplementary Materials of Angelopoulos et al. (2023a) and can be used to describe several scenarios when $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is guaranteed to have lower asymptotic variance than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when the ML model captures the true regression function (see Appendix A.4). Here we consider a simple example to demonstrate when $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ gains efficiency over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ under homoskedasticity.
Example 1 (Efficiency gain of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ under homoskedasticity)
In the homoskedastic setting where $\mathbb{E}(\epsilon_{Y}^{2}\mid{\bf{X}})=\sigma_{Y}^{2}$, $\mathbb{E}(\epsilon_{\widehat{Y}}^{2}\mid{\bf{X}})=\sigma_{\widehat{Y}}^{2}$, and $\mathbb{E}(\epsilon_{Y}\epsilon_{\widehat{Y}}\mid{\bf{X}})=\sigma_{Y,\widehat{Y}}$,
$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})=\left(\frac{1}{\pi}-1\right)\left[2\sigma_{Y,\widehat{Y}}-\sigma_{\widehat{Y}}^{2}\right]{\bf{A}},$
and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ if and only if
$\rho_{Y,\widehat{Y}}>\frac{1}{2}\frac{\sigma_{\widehat{Y}}}{\sigma_{Y}},$
where $\rho_{Y,\widehat{Y}}=\sigma_{Y,\widehat{Y}}/\sigma_{Y}\sigma_{\widehat{Y}}$ is the correlation of residuals for $Y$ and $\widehat{Y}$ given ${\bf{X}}$.
The improvement of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ is therefore stronger with increasing correlation of residuals between $Y$ and $\widehat{Y}$ and decreasing variability in the residuals of $\widehat{Y}$. Therefore, $\widehat{Y}$ must be a strong predictor for $Y$ after adjusting for ${\bf{X}}$, though it does not necessarily need to be derived from a correctly specified ML model. In particular, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ gains efficiency if and only if the correlation of residuals is at least half of the ratio of residual standard deviations of $\widehat{Y}$ to that of $Y$.
3.2 Asymptotic analysis of the CC estimator
To further analyze the efficiency of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ we draw on a landmark paper that introduced a class of augmented inverse probability weighted (AIPW) estimators which includes all regular asymptotic linear estimators (Robins et al., 1994). In the current setting, where the goal is to estimate linear regression coefficients $\bm{\beta}^{*}$ with $Y$ partially observed, but $\widehat{Y}$ fully observed, these estimators have influence functions of the form
$\text{IF}(Y,\widehat{Y},{\bf{X}};\bm{\omega})={\bf{A}}\left[\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})\frac{R}{\pi}+\bm{\omega}(\widehat{Y},{\bf{X}})\frac{\pi-R}{\pi}\right],$
with $\bm{\omega}(\widehat{Y},{\bf{X}})$ being a function of $\widehat{Y}$ and ${\bf{X}}$. The optimal choice of $\bm{\omega}(\widehat{Y},{\bf{X}})$ with respect to asymptotic variance is
$\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})=\mathbb{E}[\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})\mid\widehat{Y},{\bf{X}}].$
By Theorem 1, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is an AIPW estimator and can be seen as an approximation to the optimal estimator with $\bm{\omega}(\widehat{Y},{\bf{X}})=\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})$. The PPI estimator is therefore not the most efficient estimator in its class and, from Corollary 1.1, can even be less efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when $\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}\succ 2\boldsymbol{\Sigma}_{Y,\widehat{Y}}$. A PB inference strategy should enhance statistical efficiency relative to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ whenever possible and, at the very least, should not be negatively impacted by incorporating unlabeled data into estimation. This raises concerns about the practical utility of the PPI estimator.
To improve upon $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, one approach is to assume a parametric model for $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$ that can be estimated from $\mathscr{L}$ for analysis. However, it is often difficult to specify an appropriate parametric model for $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$, particularly when the relationship between $Y$ and $\widehat{Y}$ is unknown. Moreover, when the model for $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$ is misspecified, an efficiency gain is not guaranteed. To overcome these difficulties, Chen and Chen (2000) proposed a particular choice of $\bm{\omega}(\widehat{Y},{\bf{X}})$ that is guaranteed to be at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and that does not require modeling of $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$. Their approach is motivated by the double-sampling design which, apart from differences in terminology and notation (i.e., they regard $\widehat{Y}$ as a mismeasured version of $Y$), is identical to the partially labeled data setting considered here. The CC estimator approximates $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$ with
$\bm{\omega}^{\text{\tiny{CC}}}(\widehat{Y},{\bf{X}})=\text{Cov}[\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*}),\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})]\{\text{Var}[\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})]\}^{-1}\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*}).$ (1)
The estimator with this influence function can be written as
$\hat{\bm{\beta}}^{\text{\tiny{CC}}}=\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\widehat{{\bf{W}}}^{\text{\tiny{CC}}}\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}\right),$
where $\widehat{{\bf{W}}}^{\text{\tiny{CC}}}$ is a consistent estimator of
${{\bf{W}}}^{\text{\tiny{CC}}}={\bf{A}}{\text{Cov}}[\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*}),\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})]\{{\text{Var}}[\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})]\}^{-1}{\bf{A}}^{-1}$
obtained with $\mathscr{L}$. Therefore, a simple way to ensure that an augmented estimator like the PPI estimator has asymptotic variance at least as low as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ is to add a simple weight function in front of the augmentation term. In fact, the framework of Robins et al. (1994) implies that all regular asymptotically linear estimators for $\bm{\beta}^{*}$ augment $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ with an augmentation term that incorporates information from $\widehat{Y}$. We present the influence function expansion for $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ in Theorem 2 to further study its asymptotic variance and compare it with $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$. All of the asymptotic results provided for $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ are detailed in Appendix C.
Theorem 2 (Influence function expansion of $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$)
Following Chen and Chen (2000),
$n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{CC}}}-\bm{\beta}^{*}\right)=n^{-\frac{1}{2}}\sum_{i=1}^{n}{\bf{IF}}^{\text{\tiny{CC}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})+o_{p}(1),$
where ${\bf{IF}}^{\text{\tiny{CC}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})={\bf{A}}\left[\bm{\phi}(Y_{i},{\bf{X}}_{i};\bm{\beta}^{*})\frac{R_{i}}{\pi}+\bm{\omega}^{\text{\tiny{CC}}}(\widehat{Y}_{i},{\bf{X}}_{i})\frac{\pi-R_{i}}{\pi}\right]$ and $\bm{\omega}^{\text{\tiny{CC}}}(\widehat{Y},{\bf{X}})$ is defined in (1). Therefore, $n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{CC}}}-\bm{\beta}^{*}\right)$ converges weakly to a zero-mean multivariate normal distribution with variance-covariance
$\hskip 2.84526pt\mathbb{E}\left[{\bf{IF}}^{\text{\tiny{CC}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})\hskip 2.84526pt{\bf{IF}}^{\text{\tiny{CC}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})^{{\sf\scriptscriptstyle{T}}}\right]\hskip 2.84526pt.$
The subsequent corollaries apply Theorem 2 to show our two key results: (1) $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and (2) $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ has the highest asymptotic efficiency within the class of estimators with a weighted linear augmentation term. Together these results illustrate that the CC method is both robust to ML model specification and optimizes statistical efficiency, making it preferable to the PPI method in practice.
Corollary 2.1 ($\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$)
The difference in the asymptotic variance-covariance of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is
$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{CC}}})=\left(\frac{1}{\pi}-1\right){\bf{A}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}\boldsymbol{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}{\bf{A}}\succeq{\bf{0}}.$
Therefore, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is guaranteed to be at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
Corollary 2.2 ($\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is the most efficient estimator in its class)
Let the mean squared error (MSE) of an estimator $\hat{\bm{\beta}}$ for $\bm{\beta}^{*}$ be defined as $\mathbb{E}[(\hat{\bm{\beta}}-\bm{\beta}^{*})^{{\sf\scriptscriptstyle{T}}}(\hat{\bm{\beta}}-\bm{\beta}^{*})]$. Analogously to the well-known result for the best linear predictor (e.g., Bickel and Doksum (2015), page 40), $\widehat{{\bf{W}}}^{\text{\tiny{CC}}}$ is the weight, ${\bf{W}}$, that minimizes the asymptotic MSE of estimators of the form $\hat{\bm{\beta}}^{\text{\tiny{lab}}}-{{\bf{W}}}\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}\right)$ for estimating $\bm{\beta}^{*}$.
From Corollary 2.1, we note that in contrast to $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})$, the covariance term in $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{CC}}})$ is normalized by the residual variance of the $\widehat{Y}$ to scale the covariance according to the precision of $\widehat{Y}$. This scaling essentially enables $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ to always be as least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. Coming back to the example from the previous section, we can gain more intuition as to why $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is guaranteed to be at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ while $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ does not possess this property.
Example 1 continued (Efficiency gain of $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ under homoskedasticity) We revisit the case of homoskedasticity and recall that
$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})$ $=\left(\frac{1}{\pi}-1\right)\left(2\sigma_{Y,\widehat{Y}}-\sigma_{\widehat{Y}}^{2}\right){\bf{A}}$
$=\left(\frac{1}{\pi}-1\right)\left[-\left(\sigma_{Y}\rho_{Y,\widehat{Y}}-\sigma_{\widehat{Y}}\right)^{2}+\sigma_{Y}^{2}\rho_{Y,\widehat{Y}}^{2}\right]{\bf{A}}.$
Theorem 2 can be used to show that
$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{CC}}})$ $=\left(\frac{1}{\pi}-1\right)\left(\frac{\sigma_{Y,\widehat{Y}}^{2}}{\sigma_{\widehat{Y},\widehat{Y}}^{2}}\right){\bf{A}}=\left(\frac{1}{\pi}-1\right)\sigma_{Y}^{2}\rho_{Y,\widehat{Y}}^{2}{\bf{A}}.$
$\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ provided that $\rho_{Y,\widehat{Y}}>0$ and the gain in efficiency scales directly with $\rho_{Y,\widehat{Y}}$ and $\sigma_{Y}^{2}$. In contrast, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ will only be more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ provided that $\rho_{Y,\widehat{Y}}>\frac{1}{2}\frac{\sigma_{\widehat{Y}}}{\sigma_{Y}}$ and the gain in efficiency will only scale in this manner with the additional condition that $\rho_{Y,\widehat{Y}}=\frac{\sigma_{\widehat{Y}}}{\sigma_{Y}}$.
4 Related literature
Existing PB inference literature has established connections with semi-supervised inference, semi-parametric inference, missing data, and measurement error (Angelopoulos et al., 2023a, b; Gan and Liang, 2023; Miao et al., 2023; Zrnic and Candès, 2024a, b). Here we draw connections between PB inference methods and methods from the economics and statistics literature dating back to the 1960s. These connections reveal that aspects of the PB inference problem have been explored in various areas, albeit with differences in motivation, terminology, and notation. Given the close relationship between the PPI and CC methods, future research in PB inference could benefit from a more comprehensive review of these and other related approaches than what is provided here.
4.1 Seemingly unrelated regression & its use in statistical genetics
Seemingly unrelated regression (SUR) was introduced by economist Arnold Zellner in 1962 for simultaneous estimation of multiple related regression models (Zellner, 1962; Swamy and Mehta, 1975; Schmidt, 1977). The regression models have different outcome variables, but are assumed to have correlated errors. SUR leverages the correlation of the error terms to produce more efficient estimators compared to estimating each regression independently. In the current setting, the SUR problem can be stated formally as
$Y={\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\beta}^{*}+\epsilon\quad\mbox{and}\quad\widehat{Y}={\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma}^{*}+\epsilon^{\prime}\quad$
where $\mathbb{E}(\epsilon\mid{\bf{X}})=0$ and $\mathbb{E}(\epsilon^{\prime}\mid{\bf{X}})=0.$ An estimator for $\bm{\beta}^{*}$ that is based on augmenting $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ was proposed by Conniffe (1985) and is given by:
$\hat{\bm{\beta}}^{\text{\tiny{SUR}}}=\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\widehat{W}^{\text{\tiny{SUR}}}\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}\right)\quad\mbox{where}\quad\widehat{W}^{\text{\tiny{SUR}}}=\frac{\sum_{i=1}^{n_{\text{\tiny{lab}}}}(Y_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\hat{\bm{\beta}}^{\text{\tiny{lab}}})(\widehat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\hat{\bm{\gamma}}^{\text{\tiny{lab}}})}{\sum_{i=1}^{n_{\text{\tiny{lab}}}}(\widehat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\hat{\bm{\gamma}}^{\text{\tiny{lab}}})^{2}}.$
Note that $\widehat{W}^{\text{\tiny{SUR}}}$ is an estimator of the scalar $W^{\text{\tiny{SUR}}}=\frac{\mathbb{E}[\text{Cov}(\epsilon,\epsilon^{\prime}\mid{\bf{X}})]}{\mathbb{E}[\text{Var}(\epsilon^{\prime}\mid{\bf{X}})]}$. Similar to the PPI and CC methods, the contribution of the augmentation term to estimation increases with the predictiveness of $\widehat{Y}$ for $Y$. In the extreme case where $\widehat{Y}$ is a perfect predictor for $Y$ such that $\widehat{Y}=Y$, $\widehat{W}^{\text{\tiny{SUR}}}=1$ and $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ is simply $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$. In the homoskedastic case when $\text{Var}(\epsilon\mid{\bf{X}})$ and $\text{Cov}(\epsilon,\epsilon^{\prime}\mid{\bf{X}})$ are constant and do not depend on ${\bf{X}}$, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is asymptotically equivalent to $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ (Chen and Chen, 2000). Moreover, when the conditional distribution of $(\epsilon,\epsilon^{\prime})$ given ${\bf{X}}$ is bivariate normal with mean $\boldsymbol{0}$ and has constant variance-covariance, $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ (and thus $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ as well) is the maximum likelihood estimator of $\bm{\beta}^{*}$ and fully efficient under the assumed parametric model (Conniffe, 1985). In other scenarios, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ achieves a lower asymptotic variance than $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$.
The SUR estimator was recently applied for a PB inference problem inspired by genome-wide association studies conducted with population biobanks. These studies often rely on pre-trained models to impute outcomes of interest (i.e., phenotypes or traits) as they are partially missing due to the time and expense of ascertainment (McCaw et al., 2024). The proposed method, “SynSurr”, assumes the true and predicted outcomes follow a bivariate normal distribution. The SynSurr estimator is equivalent to $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ and asymptotically equivalent to $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ when $\text{Var}(\epsilon\mid{\bf{X}})$ and $\text{Cov}(\epsilon,\epsilon^{\prime}\mid{\bf{X}})$ are constant. Building on the SynSurr estimator, Miao et al. (2024) proposed the “POP-GWAS” estimator for a variety of regression models utilizing the augmentation technique. The POP-GWAS estimator is
$\hat{\bm{\beta}}^{\text{\tiny{POP}}}=\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\widehat{W}^{\text{\tiny{POP}}}\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}\right)\quad\mbox{where}\quad\widehat{W}^{\text{\tiny{POP}}}=\frac{n_{\text{\tiny{unlab}}}}{n}\widehat{\text{Corr}}(Y,\widehat{Y}\mid{\bf{X}}).$
In contrast to the SUR estimator, $\hat{\bm{\beta}}^{\text{\tiny{POP}}}$ uses $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ instead of $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$ for the augmentation, and the weight, $\widehat{W}^{\text{\tiny{POP}}}$ is not generally a consistent estimator of $W^{\text{SUR}}$. $\hat{\bm{\beta}}^{\text{\tiny{POP}}}$ is therefore less efficient than $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ when the residuals are bivariate normal with constant variance-covariance and less efficient than $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ in general.
4.2 Augmented estimation for quantifying a treatment effect
Outside of statistical genetics, the augmented estimation approach used by the PPI and CC methods has been widely used for treatment effect estimation in the statistics and biostatistics literature (e.g., Leon et al. (2003); Lu and Tsiatis (2008); Tsiatis et al. (2008); Gilbert et al. (2009); Zhang and Gilbert (2010); Tian et al. (2012); Yang and Ding (2020)). Many of these approaches build on methods from the survey sampling literature (e.g., Cassel et al. (1976); Särndal et al. (2003)).
For example, Tsiatis et al. (2008) consider covariate adjusted analyses for comparing two treatments in randomized clinical trials. Letting $X$ be an indicator for randomization to the treated or control arms, the focus of inference is on $\beta$ in the model
$\mathbb{E}(Y\mid X)=\alpha+\beta X$
in order to estimate the treatment effect $\Delta=\mathbb{E}(Y\mid X=1)-\mathbb{E}(Y\mid X=0)$. A standard estimator of $\Delta$ is $\widehat{\Delta}=\overline{Y}^{1}-\overline{Y}^{0}$, where $\overline{Y}^{1}=n_{1}^{-1}\sum_{i=1}^{n}Y_{i}X_{i}$ and $\overline{Y}^{0}=n_{0}^{-1}\sum_{i=1}^{n}Y_{i}(1-X_{i})$ are the empirical means of $Y$ in the treated and control groups of sizes $n_{1}$ and $n_{0}$, respectively. To utilize information in baseline covariates, ${\bf{Z}}$, Tsiatis et al. (2008) propose an augmented estimator of the treatment effect given by
$\widehat{\Delta}-\sum_{i=1}^{n}(X_{i}-\bar{X})[n_{0}^{-1}h^{(0)}({\bf{Z}}_{i})+n_{1}^{-1}h^{(1)}({\bf{Z}}_{i})],$ (2)
where $\bar{X}=n^{-1}\sum_{i=1}^{n}X_{i}$ and $h^{(k)}({\bf{Z}})$ for $k=0,1$ are arbitrary scalar functions of ${\bf{Z}}$. As $X\perp{\bf{Z}}$ are independent due to randomization, the augmentation term in 2 converges in probability to 0, analogously to the augmentation terms in the PPI and CC estimators. In the context of PB inference, pre-trained models could be used for predicting $Y$ on the basis of ${\bf{Z}}$ for $h^{(k)}({\bf{Z}})$ for $k=0,1$.
4.3 Improving precision of Monte Carlo simulations with control variates
The augmentation technique has also been utilized to improve the efficiency of Monte Carlo estimation. Specifically, Rothery (1982) considers using a control variate, a statistic whose properties are known, to improve precision when estimating power via simulation. For exposition, suppose interest lies in estimating the power of a test $T$ of $H_{0}:\theta=0$ for some scalar parameter $\theta$. Let $S$ denote an existing test of the same hypothesis whose power $\mu_{Z}=\mathbb{E}_{\theta}(Z)$ is known. Define $Y$ as an indicator that test $T$ rejects and $Z$ as an indicator that test $S$ rejects. For a set of $n$ simulation replicates, the baseline estimator of power for $T$ is simply $\overline{Y}=n^{-1}\sum_{i=1}^{n}Y_{i}$ while a control variate augmented estimator is given by:
$\frac{1}{n}\sum_{i=1}^{n}\left[Y_{i}-\lambda\{Z_{i}-\mu_{Z}\}\right].$ (3)
The variance of the augmented estimator is minimized at $\lambda^{*}=\frac{\text{Cov}(Y,Z)}{\text{Var}(Z)}$ Rothery (1982); Davidson and MacKinnon (1992). The augmentation term reduces the variance of the unadjusted power estimate $\overline{Y}$ in proportion to the correlation between $Y$ and $Z$. Building on work from the survey sampling literature, a similar idea was later used for mean estimation in the semi-supervised setting by Zhang et al. (2019).
5 Numerical studies
5.1 Data generation
We perform several simulation studies to compare the performance of the PPI and CC methods. In all settings, we generate two datasets: (i) a training dataset with realizations of $(Y,{\bf{Z}}^{{\sf\scriptscriptstyle{T}}},X_{1})^{{\sf\scriptscriptstyle{T}}}$ to train the ML model and (ii) an independent analysis dataset for evaluating the PB inference methods containing a labeled dataset with realizations of $(Y,{\bf{Z}}^{{\sf\scriptscriptstyle{T}}},X_{1})^{{\sf\scriptscriptstyle{T}}}$ and an unlabeled dataset with realizations of $({\bf{Z}}^{{\sf\scriptscriptstyle{T}}},X_{1})^{{\sf\scriptscriptstyle{T}}}$. We use 10,000 replications of the analysis dataset for each scenario and a single training set as $\hat{f}$ is assumed to be fixed. The size of the training set is fixed at 10,000 and the size of the analysis dataset is also set to $n=10,000$. We let $n_{\text{\tiny{lab}}}=1000,2000,3000,4000,5000$ and $n_{\text{\tiny{unlab}}}=n-n_{\text{\tiny{lab}}}$. For both the training and analysis datasets, $({\bf{Z}}^{{\sf\scriptscriptstyle{T}}},X_{1})^{{\sf\scriptscriptstyle{T}}}$ are generated from a multivariate normal distribution. Each element of $({\bf{Z}}^{{\sf\scriptscriptstyle{T}}},X_{1})^{{\sf\scriptscriptstyle{T}}}$ has a mean of 0, variance of 1, and covariance of 0.4 with the remaining variables. The goal is to conduct inference on the marginal association of $Y$ and $X_{1}$ in a linear model. The parameter of interest is therefore $\beta^{*}_{1}$ defined as
$(\beta^{*}_{0},\beta^{*}_{1})=\operatorname*{arg\,min}_{\bm{\beta}}\mathbb{E}[(Y-(\beta_{0}+\beta_{1}X_{1}))^{2}].$
We consider 2 primary scenarios for evaluating the PB inference methods. In both scenarios, the outcome is generated as
$Y=X_{1}+\theta_{\text{\tiny{INT}}}X_{1}W_{1}+{\bf{W}}^{{\sf\scriptscriptstyle{T}}}\bm{\theta}+\epsilon$
where ${\bf{W}}^{{\sf\scriptscriptstyle{T}}}=(W_{1},\dots,W_{q})$ and the error term $\epsilon$ is simulated independently from a standard normal distribution. We varied the coefficients of $\bm{\theta}$ and $\theta_{\text{\tiny{INT}}}$ as well as $q$ and the composition of ${\bf{Z}}$.
Scenario 1
We consider two sub-scenarios to illustrate Corollary 1.1, that is, that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when the sufficient condition holds. In both sub-scenarios, we let $p=10$, $\bm{\theta}=(1,1,0.5,0.5,0.25,{\bf{0}}_{1\times 5})^{{\sf\scriptscriptstyle{T}}}$, and $\theta_{\text{\tiny{INT}}}=0$. In Scenario 1a, we obtain a strong prediction model by regressing $Y$ on ${\bf{Z}}=(X_{1},W_{1},\dots,W_{10})^{{\sf\scriptscriptstyle{T}}}$ while in Scenario 1b we obtain a weaker prediction model by regressing $Y$ on ${\bf{Z}}=(X_{1},W_{5},X_{1}W_{5})^{{\sf\scriptscriptstyle{T}}}$.
Scenario 2
We consider two sub-scenarios to illustrate the results in Corollary 1.1 and Corollary 2.1, which show that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ can be less efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ with a weak prediction model while $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is at least as efficient $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. In both scenarios, we mimic a common scenario where the prediction model does not contain $X_{1}$ (e.g., in a genome-wide association study). In Scenario 2a, we let $p=9$, $\bm{\theta}=(0.2,0.1\times{\bf{1}}_{1\times 8})^{{\sf\scriptscriptstyle{T}}}$, and $\theta_{\text{\tiny{INT}}}=0$. We obtain the prediction model by regressing $Y$ on ${\bf{Z}}=(W_{1},\dots,W_{9})^{{\sf\scriptscriptstyle{T}}}$. In Scenario 2b, we let $p=3$, $\bm{\theta}=(0.25,0.25,0.25)^{{\sf\scriptscriptstyle{T}}}$, and $\theta_{\text{\tiny{INT}}}=0.25$. We obtain the prediction model by regressing $Y$ on ${\bf{Z}}=(W_{1},W_{2},W_{3},W_{2}W_{1},W_{3}W_{1},W_{3}W_{2})^{{\sf\scriptscriptstyle{T}}}$. Scenario 2b uses a weaker prediction model than Scenario 2a.
We evaluate the performance of the PPI and CC methods: $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, and $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$. We also include two comparison methods: $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, serving as the baseline, and $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$, highlighting the limitations of naive inference. We present the percent bias for each estimator. For the PPI and CC methods, we present the confidence interval (CI) length ratio, computed as the mean of the ratio of the length of the PB inference estimator’s normal-based CI relative to the length of the normal-based CI for $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ (which is simply the ratio of the asymptotic standard errors). The CI length ratio has a baseline value of 1, which indicates that the CI of the PB inference method is the same length as the CI for $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. A CI length ratio less than 1 indicates that the PB inference estimator has lower standard error than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
5.2 Results
Table 1 presents the percent bias of the various estimators across the different scenarios and labeled dataset sizes. With the exception of $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$, the methods exhibit minimal bias as they are all consistent estimators for $\bm{\beta}^{*}_{1}$. The limitations of naive inference are particularly apparent in Scenarios 2a and 2b, which do not include $X_{1}$ in ML model training. $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$ exhibits an average percent bias of -47% in Scenario 2a and -64% in Scenario 2b.
Figure 2 displays the CI length ratio across the different settings. In Scenario 1, we observe that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ and $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ perform similarly, with CI lengths much less 1, indicating that the PB inference methods have reduced variance relative to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. Comparing Scenario 1a and 1b, we note a significant decrease in the gains in statistical efficiency of the PB inference methods due to the weaker prediction model used in Scenario 1b. However, all methods are at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, as the sufficient condition in Corollary 1.1 is satisfied. Additionally, as the labeled set size increases and hence $\pi=n_{\text{\tiny{lab}}}/n$ increases, the efficiency gain decreases as expected. $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is always more efficient than $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, particularly when the labeled dataset size is larger, due to the difference in the augmentation terms of these two estimators. In this scenario, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is asymptotically equivalent to $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$.
Scenario 2 highlights the differences in performance between $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ and $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$, which underscores our findings in Corollary 1.1 and Corollary 2.1. In both sub-scenarios, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ consistently outperforms both $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ and is at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. In Scenario 2a, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ achieves an average CI length ratio of 0.9 across the various sample sizes, indicating a gain in efficiency relative to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. In Scenario 2b, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is comparable to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ for the larger labeled dataset sizes. In contrast, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ are outperformed by $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and have confidence intervals up to 30% longer than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ for the larger labeled dataset sizes. This behavior is exacerbated in Scenario 2b, where the CI of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is up to 50% longer than the CI for $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and up to 72% longer for $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$.
Labeled set size
Method 1000 2000 3000 4000 5000
Scenario 1a
     $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$0.181 0.124 0.650 0.430 0.165
     $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ 0.035 -0.042 0.015 0.005 -0.005
     $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ 0.004 0.010 -0.004 -0.018 -0.001
     $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ 0.007 0.000 0.002 -0.009 -0.003
     $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ 0.007 0.000 0.002 -0.009 -0.003
Scenario 1b
     $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$1.070 0.181 0.426 -0.423 -1.205
     $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ 0.008 0.002 -0.016 -0.001 0.004
     $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ 0.007 -0.001 -0.003 -0.008 -0.010
     $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ 0.007 $0$0.007 -0.006 -0.003
     $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ 0.006 $0$0.007 -0.006 -0.003
Scenario 2a
     $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$46.980 -47.523 -45.946 -46.725 -47.695
     $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$0.033 -0.023 0.027 0.004 0.005
     $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$0.023 -0.009 -0.003 -0.032 -0.003
     $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$0.024 -0.012 0.006 -0.018 0.001
     $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$0.030 -0.016 0.016 -0.007 0.003
Scenario 2b
     $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$64.272 -64.115 -63.307 -63.759 -64.627
     $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$0.012 -0.003 0.002 0.012 0.013
     $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ 0.000 0.001 -0.006 0.012 0.009
     $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$0.001 $0$0.004 0.012 0.010
     $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$0.014 -0.006 -0.003 0.011 0.010
Table 1: Percent biases of the various estimators. Negative values indicate underestimation, while positive values indicate overestimation of the true parameter.
Figure 2: Confidence interval length ratio of the PB inference methods compared to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. A ratio less than 1 indicates that the PB inference method has a confidence interval that is shorter than the confidence interval for $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
6 Discussion
In this paper, we showed that adding a simple weight to the PPI estimator provides valid inference and is guaranteed to be at least as efficient as estimation using only labeled data. The weight is drawn from the work of Chen and Chen (2000) in the context of double sampling. In contrast, the original PPI estimator can be less efficient than estimation using only the labeled data, a property that can limit its practical utility. By incorporating the weight derived from Chen and Chen (2000), the augmented estimation approach principally balances the contribution of labeled and unlabeled data and extends the applicability of the PPI method to scenarios where leveraging all available data is essential for making valid and efficient inference. In addition to theoretical results that support these findings, our simulation studies demonstrate that the PPI method can result in substantial efficiency loss when the ML model is inaccurate. Additionally, we provide connections to several existing approaches from the economics and statistics literature that provide some interesting context for the PB inference problem. While we focused on least squares estimation, these results extend to other settings. The work of Chen and Chen (2000) and Angelopoulos et al. (2023a) apply to a wide variety of models beyond linear regression, such as parameters defining generalized linear models or parameters derived from estimating equations.
The work of Angelopoulos et al. (2023a) was pivotal in advancing the growing body of literature in PB inference. Recent work has considered settings where the ML model is not black box and where active learning is utilized to obtain the labeled data Zrnic and Candès (2024b, b). Additionally, two concurrent pre-prints have noted potential losses in efficiency of PB inference approaches. The authors of the PPI paper introduced a new method, PPI++, which aims to address this limitation of the PPI estimator (Angelopoulos et al., 2023b). The PPI++ method is designed to adaptively minimize the trace of the asymptotic variance-covariance of the proposed estimator to ensure that it yields an improvement over an approach based solely on labeled data. In contrast, the approach of Chen and Chen (2000) minimizes the mean square error among a class of estimators with a linear augmentation term. Simultaneously, Miao et al. (2023) introduced the PoSt-Prediction Adaptive inference (PSPA) estimator that is both robust to the accuracy of the ML model and guaranteed to be at least as efficient as estimation using only labeled data. Unlike the PPI++ estimator, the PSPA estimator is designed to achieve element-wise variance reduction and was shown to outperform the PPI++ approach. The PSPA approach is based on augmenting estimating equations with information from $\widehat{Y}$, while the CC method provides a simple and natural extension to the PPI method by adding a weight to the augmentation term.
Code availability
Our analyses can be replicated with the code at: https://github.com/SelinaS37/PBInference.
Acknowledgments
The authors gratefully acknowledge funding from an NSERC Discovery Grant (RGPIN-2021-03734), the Connaught Fund, the University of Toronto Data Science Institute, and the McLaughlin Center for Genetic Research.
Appendix A Asymptotic properties of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$
A.1 Preliminaries
Notation. We denote the outer product of a vector $\boldsymbol{v}$ as $\boldsymbol{v}^{\otimes 2}=\boldsymbol{v}\boldsymbol{v}^{{\sf\scriptscriptstyle{T}}}$. We denote that a matrix, ${\bf{M}}$, is positive semi-definite as ${\bf{M}}\succeq{\bf{0}},$ and positive definite as ${\bf{M}}\succ{\bf{0}}.$ For any estimator $\hat{\bm{\beta}}$, we let $\text{aVar}(\hat{\bm{\beta}})$ represent its asymptotic variance-covariance matrix. We define the difference in asymptotic variance-covariance matrices between two estimators $\hat{\bm{\beta}}_{1}$ and $\hat{\bm{\beta}}_{2}$ as $\Delta_{\text{aVar}}(\hat{\bm{\beta}}_{1},\hat{\bm{\beta}}_{2})=n[\text{aVar}(\hat{\bm{\beta}}_{1})-\text{aVar}(\hat{\bm{\beta}}_{2})].$
Basic assumptions. We make the standard assumption that $\text{Var}({\bf{X}})\succ{\bf{0}}$. We also assume that ${\bf{X}}$ has compact support and that $(Y_{i},\widehat{Y}_{i},{\bf{X}}^{{\sf\scriptscriptstyle{T}}})^{{\sf\scriptscriptstyle{T}}}$ have finite first and second moments.
A.2 Proof of Theorem 1
Theorem 1 (Influence function expansion of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$)
Let $R_{i}$ be the indicator of whether the $i$th observation is labeled. Then,
$n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}-\bm{\beta}^{*}\right)=n^{-\frac{1}{2}}\sum_{i=1}^{n}{\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})+o_{p}(1)$
where ${\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})={\bf{A}}\left\{\left[\ \bm{\phi}(Y_{i},{\bf{X}}_{i};\bm{\beta}^{*})\right]\frac{R_{i}}{\pi}+\left[\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma}^{*})\right]\frac{\pi-R_{i}}{\pi}\right\}$ and ${\bf{A}}=[\mathbb{E}({\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}})]^{-1}$. Therefore, $n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}-\bm{\beta}^{*}\right)$ converges weakly to a zero-mean multivariate normal distribution with variance-covariance
$\mathbb{E}[{\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})\hskip 2.84526pt{\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})^{{\sf\scriptscriptstyle{T}}}].$
Proof  First, note that $n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}-\bm{\beta}^{*}\right)$ may be decomposed as
$n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}-\bm{\beta}^{*}\right)=n^{\frac{1}{2}}\left[\left(\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\bm{\beta}^{*}\right)-\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\bm{\gamma}^{*}\right)+\left(\hat{\bm{\gamma}}^{\text{\tiny{all}}}-\bm{\gamma}^{*}\right)\right].$
We consider the asymptotic linear expansions of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, $\hat{\bm{\gamma}}^{\text{\tiny{lab}}}$, and $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$. Since $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ is the familiar least squares estimator,
$n_{\text{\tiny{lab}}}^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\bm{\beta}^{*}\right)=\left(n_{\text{\tiny{lab}}}^{-1}\sum_{i=1}^{n_{\text{\tiny{lab}}}}{\bf{X}}_{i}{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\right)^{-1}n_{\text{\tiny{lab}}}^{-\frac{1}{2}}\sum_{i=1}^{n_{\text{\tiny{lab}}}}{\bf{X}}_{i}(Y_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\bm{\beta}^{*}).$ (4)
By Slutsky’s theorem and the assumption that $\lim_{n\to\infty}n_{\text{\tiny{lab}}}/n=\pi$, equation (4) can be rewritten as
$\begin{split}&n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\bm{\beta}^{*}\right)=n^{-\frac{1}{2}}\sum_{i=1}^{n}{\bf{A}}\left\{\left[{\bf{X}}_{i}\left(Y_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\bm{\beta}^{*}\right)\right]\frac{R_{i}}{\pi}\right\}+o_{p}(1).\end{split}$ (5)
Similar arguments can also be used to show that
$n^{\frac{1}{2}}\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\bm{\gamma}^{*}\right)=n^{-\frac{1}{2}}\sum_{i=1}^{n}{\bf{A}}\left\{\left[{\bf{X}}_{i}\left(\widehat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma}^{*}\right)\right]\frac{R_{i}}{\pi}\right\}+o_{p}(1)$
and
$n^{\frac{1}{2}}\left(\hat{\bm{\gamma}}^{\text{\tiny{all}}}-\bm{\gamma}^{*}\right)=n^{-\frac{1}{2}}\sum_{i=1}^{n}{\bf{A}}\left\{\left[{\bf{X}}_{i}\left(\widehat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma}^{*}\right)\right]\right\}+o_{p}(1).$
Consequently,
$n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}-\bm{\beta}^{*}\right)$ $=n^{-\frac{1}{2}}\sum_{i=1}^{n}{\bf{A}}\left\{\left[\bm{\phi}(Y_{i},{\bf{X}}_{i};\bm{\beta}^{*})\frac{R_{i}}{\pi}+\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i},\bm{\gamma}^{*})\frac{\pi-R_{i}}{\pi}\right]\right\}+o_{p}(1)$
$=n^{-\frac{1}{2}}\sum_{i=1}^{n}{\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})+o_{p}(1).$
It follows from the standard central limit theorem (CLT) that $n^{\frac{1}{2}}(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}-\bm{\beta}^{*})$ converges weakly to a zero mean multivariate Gaussian distribution with variance-covariance given by $\mathbb{E}\left[{\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})\hskip 2.84526pt{\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})^{{\sf\scriptscriptstyle{T}}}\right].$
A.3 Proof of Corollary 1.1
Corollary 1.1 (Comparison of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ with $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$)
The difference in the asymptotic variance-covariance of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is
$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})=\left(\frac{1}{\pi}-1\right){\bf{A}}\left(2\boldsymbol{\Sigma}_{Y,\widehat{Y}}-\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}\right){\bf{A}}.$
$\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is therefore more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ if and only if $2\boldsymbol{\Sigma}_{Y,\widehat{Y}}\succ\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}$. A sufficient condition for $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ to have lower variance-covariance than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ is that
$\mathbb{E}\left[\left(\epsilon_{\widehat{Y}}-\epsilon_{Y}\right)^{2}\mid{\bf{X}}\right]<\mathbb{E}\left(\epsilon_{Y}^{2}\mid{\bf{X}}\right).$
Proof  Equation (5) in the proof of Theorem 1 gives the asymptotic expansion of $n^{\frac{1}{2}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\bm{\beta})$. The standard CLT implies $n^{\frac{1}{2}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\bm{\beta})$ converges weakly to a zero-mean multivariate Gaussian distribution with variance-covariance given by
$\mathbb{E}\left\{\left[{\bf{A}}\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})\right]^{\otimes 2}\frac{R^{2}}{\pi^{2}}\right\}=\frac{1}{\pi}{\bf{A}}\boldsymbol{\Sigma}_{Y,Y}{\bf{A}}.$
Similarly, by Theorem 1
$\begin{split}n[\text{aVar}(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})]=&\mathbb{E}\left\{\left[{\bf{A}}\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})\right]^{\otimes 2}\frac{R^{2}}{\pi^{2}}\right\}+\mathbb{E}\left\{\left[{\bf{A}}\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})\right]^{\otimes 2}\frac{(\pi-R)^{2}}{\pi^{2}}\right\}\\ +&2\mathbb{E}\left\{\left[{\bf{A}}\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})\right]\left[{\bf{A}}\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})\right]^{{\sf\scriptscriptstyle{T}}}\frac{R(\pi-R)}{\pi^{2}}\right\}\\ =&\frac{1}{\pi}{\bf{A}}\left[\boldsymbol{\Sigma}_{Y,Y}+(1-\pi)\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}-2(1-\pi)\boldsymbol{\Sigma}_{Y,\widehat{Y}}\right]{\bf{A}}.\end{split}$
Thus,
$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})=\left(\frac{1}{\pi}-1\right){\bf{A}}\left(2\boldsymbol{\Sigma}_{Y,\widehat{Y}}-\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}\right){\bf{A}}.$
$\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is therefore more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ if and only if $2\boldsymbol{\Sigma}_{Y,\widehat{Y}}\succ\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}$. Additionally, if $2\mathbb{E}\left(\epsilon_{Y}\epsilon_{\widehat{Y}}\mid{\bf{X}}\right)>\mathbb{E}\left(\epsilon_{\widehat{Y}}^{2}\mid{\bf{X}}\right)$ or equivalently if $\mathbb{E}\left[\left(\epsilon_{\widehat{Y}}-\epsilon_{Y}\right)^{2}\mid{\bf{X}}\right]<\mathbb{E}\left(\epsilon_{Y}^{2}\mid{\bf{X}}\right),$ the diagonal entries of $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})$ are positive and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ will have a lower variance than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
A.4 Examples for Corollary 1.1
We provide several scenarios where the sufficient condition from Corollary 1.1 is satisfied and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. These scenarios characterize settings where the predictions contain more information about $Y$ beyond what is explained by the linear regression model based on ${\bf{X}}$ (i.e., the focus of statistical inference.)
Examples of Corollary 1.1. Let ${\bf{Z}}=({\bf{X}}^{{\sf\scriptscriptstyle{T}}},{\bf{W}}^{{\sf\scriptscriptstyle{T}}})^{{\sf\scriptscriptstyle{T}}}$ where ${\bf{W}}$ are predictors of $Y$ beyond ${\bf{X}}$. $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ in the following scenarios where $\hat{f}$ perfectly estimates the true regression function:
  1. 1.
    $\hat{f}({\bf{X}})=\mathbb{E}(Y\mid{\bf{X}})$ and $\mathbb{E}(Y\mid{\bf{X}})\neq{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\beta}^{*}$
  2. 2.
    $\hat{f}({\bf{Z}})=\mathbb{E}(Y\mid{\bf{Z}})$, ${\bf{W}}\perp Y\mid{\bf{X}}$, and $\mathbb{E}(Y\mid{\bf{X}})\neq{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\beta}^{*}$
  3. 3.
    $\hat{f}({\bf{Z}})=\mathbb{E}(Y\mid{\bf{Z}})$ and ${\bf{W}}\not\perp Y\mid{\bf{X}}$
  4. 4.
    $\hat{f}({\bf{W}})=\mathbb{E}(Y\mid{\bf{W}})$ and ${\bf{X}}\perp Y\mid{\bf{W}}$
Proof  For scenario 1, assume that $\hat{f}({\bf{X}})=\mathbb{E}(Y\mid{\bf{X}})$. It follows that
$\bm{\gamma}^{*}=\mathbb{E}({\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}})\mathbb{E}[{\bf{X}}\mathbb{E}(Y\mid{\bf{X}})]=\bm{\beta}^{*}$
so that
$\mathbb{E}\left[\left(\epsilon_{\widehat{Y}}-\epsilon_{Y}\right)^{2}\mid{\bf{X}}\right]$ $=\mathbb{E}\left\{[\mathbb{E}(Y\mid{\bf{X}})-Y]^{2}\mid{\bf{X}}\right\}$ $<\mathbb{E}\left(\epsilon_{Y}^{2}\mid{\bf{X}}\right)=\mathbb{E}\left[(Y-{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\beta}^{*})^{2}\mid{\bf{X}}\right].$
This inequality is strict provided that $\mathbb{E}(Y\mid{\bf{X}})\neq{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\beta}^{*}$. If the regression model is correctly specified (i.e., $\mathbb{E}(Y\mid{\bf{X}})={\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\beta}^{*}$), then $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ are asymptotically equivalent.
For scenario 2, assume that $\hat{f}({\bf{Z}})=\mathbb{E}(Y\mid{\bf{Z}})$ where ${\bf{Z}}=({\bf{X}}^{{\sf\scriptscriptstyle{T}}},{\bf{W}}^{{\sf\scriptscriptstyle{T}}})^{{\sf\scriptscriptstyle{T}}}$ and ${\bf{W}}\perp Y\mid{\bf{X}}$. Similar to scenario 1,
$\bm{\gamma}^{*}=\mathbb{E}({\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}})\mathbb{E}[{\bf{X}}\mathbb{E}(Y\mid{\bf{X}},{\bf{W}})]=\mathbb{E}({\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}})\mathbb{E}[{\bf{X}}\mathbb{E}(Y\mid{\bf{X}})]=\bm{\beta}^{*}.$
Therefore,
$\mathbb{E}\left[\left(\epsilon_{\widehat{Y}}-\epsilon_{Y}\right)^{2}\mid{\bf{X}}\right]$ $=\mathbb{E}\left\{[\mathbb{E}(Y\mid{\bf{Z}})-Y]^{2}\mid{\bf{X}}\right\}$
$=\mathbb{E}\left\{[\mathbb{E}(Y\mid{\bf{X}})-Y]^{2}\mid{\bf{X}}\right\}=\text{Var}(Y\mid{\bf{X}})\leq\mathbb{E}\left(\epsilon_{Y}^{2}\mid{\bf{X}}\right).$
The inequality is strict when $\mathbb{E}(Y\mid{\bf{X}})\neq{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\beta}^{*}$ and is an equality if the regression model is correctly specified.
For scenario 3, again assume that $\hat{f}({\bf{Z}})=\mathbb{E}(Y\mid{\bf{Z}})$ so that $\bm{\gamma}^{*}=\bm{\beta}^{*}$. It then follows that
$\mathbb{E}\left[\left(\epsilon_{\widehat{Y}}-\epsilon_{Y}\right)^{2}\mid{\bf{X}}\right]$ $=\mathbb{E}\left\{[\mathbb{E}(Y\mid{\bf{Z}})-Y]^{2}\mid{\bf{X}}\right\}$
$=\text{Var}(Y\mid{\bf{X}})-\text{Var}[\mathbb{E}(Y\mid{\bf{X}},{\bf{W}})\mid{\bf{X}}]$
$<\text{Var}(Y\mid{\bf{X}})\leq\mathbb{E}\left(\epsilon_{Y}^{2}\mid{\bf{X}}\right).$
The equality is strict as we assume ${\bf{W}}\not\perp Y\mid{\bf{X}}$.
Lastly, for scenario 4, assume that $\hat{f}({\bf{W}})=E(Y\mid{\bf{W}})$ and ${\bf{X}}\perp Y\mid{\bf{W}}$. Then,
$\bm{\gamma}^{*}=\mathbb{E}({\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}})\mathbb{E}[{\bf{X}}\mathbb{E}(Y\mid{\bf{W}})]=\mathbb{E}({\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}})\mathbb{E}[{\bf{X}}\mathbb{E}(Y\mid{\bf{W}},{\bf{X}})]=\bm{\beta}^{*}.$
Analogously to the argument for scenario 3,
$\mathbb{E}\left[\left(\epsilon_{\widehat{Y}}-\epsilon_{Y}\right)^{2}\mid{\bf{X}}\right]$ $=\mathbb{E}\left\{[\mathbb{E}(Y\mid{\bf{W}})-Y]^{2}\mid{\bf{X}}\right\}$
$=\text{Var}(Y\mid{\bf{X}})-\text{Var}[\mathbb{E}(Y\mid{\bf{W}})\mid{\bf{X}}]$
$<\text{Var}(Y\mid{\bf{X}})\leq\mathbb{E}\left(\epsilon_{Y}^{2}\mid{\bf{X}}\right).$
Appendix B Asymptotic properties of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$
B.1 Asymptotic expansion of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$
Theorem S1 (Influence function expansion of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$)
Let $R_{i}$ be the indicator of whether the $i$th observation is labeled. Then,
$n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{PPI}}}-\bm{\beta}^{*}\right)=n^{-\frac{1}{2}}\sum_{i=1}^{n}{\bf{IF}}^{\text{\tiny{PPI}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})+o_{p}(1)$
where ${\bf{IF}}^{\text{\tiny{PPI}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})={\bf{A}}\left\{\left[\bm{\phi}(Y_{i},{\bf{X}}_{i};\bm{\beta}^{*})\frac{R_{i}}{\pi}+\bm{\omega}^{\text{\tiny{PPI}}}(\widehat{Y}_{i},{\bf{X}}_{i},\bm{\gamma}^{*})\frac{\pi-R_{i}}{\pi}\right]\right\}$, and
$\bm{\omega}^{\text{\tiny{PPI}}}(\widehat{Y}_{i},{\bf{X}}_{i},\bm{\gamma}^{*})=\frac{\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma}^{*})}{1-\pi}$. Therefore, $n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{PPI}}}-\bm{\beta}^{*}\right)$ converges weakly to a zero-mean multivariate normal distribution with variance-covariance
$\mathbb{E}\left[{\bf{IF}}^{\text{\tiny{PPI}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})\hskip 2.84526pt{\bf{IF}}^{\text{\tiny{PPI}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})^{{\sf\scriptscriptstyle{T}}}\right].$
Proof  The only difference in studying the asymptotic properties of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ compared to $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is that we rely on the asymptotic linear expansion of $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ rather than $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$. Using the same argument as in the proof of Theorem 1,
$\begin{split}n^{\frac{1}{2}}\left(\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}-\bm{\gamma}^{*}\right)&=n^{-\frac{1}{2}}\sum_{i=1}^{n}{\bf{A}}\left\{\left[{\bf{X}}_{i}\left(\widehat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma}^{*}\right)\right]\frac{1-R_{i}}{1-\pi}\right\}+o_{p}(1).\end{split}$
Consequently,
$n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{PPI}}}-\bm{\beta}^{*}\right)$ $=n^{-\frac{1}{2}}\sum_{i=1}^{n}{\bf{A}}\left\{\left[\bm{\phi}(Y_{i},{\bf{X}}_{i};\bm{\beta}^{*})\frac{R_{i}}{\pi}+\frac{\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma}^{*})}{1-\pi}\frac{\pi-R_{i}}{\pi}\right]\right\}+o_{p}(1)$
$=n^{-\frac{1}{2}}\sum_{i=1}^{n}{\bf{IF}}^{\text{\tiny{PPI}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})+o_{p}(1).$
The standard CLT implies $n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{PPI}}}-\bm{\beta}^{*}\right)$ converges weakly to a zero-mean multivariate normal with variance-covariance $\mathbb{E}\left[{\bf{IF}}^{\text{\tiny{PPI}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})\hskip 2.84526pt{\bf{IF}}^{\text{\tiny{PPI}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})^{{\sf\scriptscriptstyle{T}}}\right].$
B.2 Comparison of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$
Corollary S1.1 (Comparison of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ with $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$)
The difference in the asymptotic variance-covariance of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is
$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}})$ $=\frac{1}{\pi}{\bf{A}}\left(2\boldsymbol{\Sigma}_{Y,\widehat{Y}}-\frac{1}{1-\pi}\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}\right){\bf{A}}.$
$\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is therefore more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ if and only if $2\boldsymbol{\Sigma}_{Y,\widehat{Y}}\succ\frac{1}{1-\pi}\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}$. A sufficient condition for $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ to have lower variance than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ is that
$\mathbb{E}\left[\left(\epsilon_{\widehat{Y}}-\epsilon_{Y}\right)^{2}\mid{\bf{X}}\right]<\mathbb{E}\left(\epsilon_{Y}^{2}\mid{\bf{X}}\right)+\left(1-\frac{1}{1-\pi}\right)\mathbb{E}\left(\epsilon_{\widehat{Y}}^{2}\mid{\bf{X}}\right).$
Proof  From the proof for Corollary 1.1, $n[\text{aVar}(\hat{\bm{\beta}}^{\text{\tiny{lab}}})]=\frac{1}{\pi}{\bf{A}}\boldsymbol{\Sigma}_{Y,Y}{\bf{A}}.$ By Theorem S1,
$n[\text{aVar}(\hat{\bm{\beta}}^{\text{\tiny{PPI}}})]=$ $\mathbb{E}\left\{\left[{\bf{A}}\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})\right]^{\otimes 2}\frac{R^{2}}{\pi^{2}}\right\}+\mathbb{E}\left\{\left[{\bf{A}}\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})\right]^{\otimes 2}\frac{(\pi-R)^{2}}{\pi^{2}(1-\pi)^{2}}\right\}$
$+2\mathbb{E}\left\{\left[{\bf{A}}\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})\right]\left[{\bf{A}}\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})\right]^{{\sf\scriptscriptstyle{T}}}\frac{R(\pi-R)}{\pi^{2}(1-\pi)}\right\}$
$=\frac{1}{\pi}{\bf{A}}\boldsymbol{\Sigma}_{Y,Y}{\bf{A}}+\frac{1}{\pi(1-\pi)}{\bf{A}}\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}{\bf{A}}-\frac{2}{\pi}{\bf{A}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}{\bf{A}}$
$=\frac{1}{\pi}{\bf{A}}\left[\boldsymbol{\Sigma}_{Y,Y}+\frac{1}{1-\pi}\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}-2\boldsymbol{\Sigma}_{Y,\widehat{Y}}\right]{\bf{A}}.$
Thus,
$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}})=\frac{1}{\pi}{\bf{A}}\left[2\boldsymbol{\Sigma}_{Y,\widehat{Y}}-\frac{1}{1-\pi}\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}\right]{\bf{A}}.$
$\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is therefore more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ if and only if $2\boldsymbol{\Sigma}_{Y,\widehat{Y}}\succ\frac{1}{1-\pi}\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}$. Note that if $2\mathbb{E}\left(\epsilon_{Y}\epsilon_{\widehat{Y}}\mid{\bf{X}}\right)>\frac{1}{1-\pi}\mathbb{E}\left(\epsilon_{\widehat{Y}}^{2}\mid{\bf{X}}\right)$ or equivalently if $\mathbb{E}\left[\left(\epsilon_{\widehat{Y}}-\epsilon_{Y}\right)^{2}\mid{\bf{X}}\right]<\mathbb{E}\left(\epsilon_{Y}^{2}\mid{\bf{X}}\right)+\left(1-\frac{1}{1-\pi}\right)\mathbb{E}\left(\epsilon_{\widehat{Y}}^{2}\mid{\bf{X}}\right),$ the diagonal entries of $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}})$ are positive and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ will have a lower variance than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
B.3 Comparison of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$
Recall from Theorem 1 and S1 that
$n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}-\bm{\beta}^{*}\right)=n^{-\frac{1}{2}}\sum_{i=1}^{n}{\bf{A}}\left\{\left[\bm{\phi}(Y_{i},{\bf{X}}_{i};\bm{\beta}^{*})\frac{R_{i}}{\pi}+\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i},\bm{\gamma}^{*})\frac{\pi-R_{i}}{\pi}\right]\right\}+o_{p}(1)\quad\mbox{and}$
$n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{PPI}}}-\bm{\beta}^{*}\right)=n^{-\frac{1}{2}}\sum_{i=1}^{n}{\bf{A}}\left\{\left[\bm{\phi}(Y_{i},{\bf{X}}_{i};\bm{\beta}^{*})\frac{R_{i}}{\pi}+\bm{\omega}^{\text{\tiny{PPI}}}(\widehat{Y}_{i},{\bf{X}}_{i},\bm{\gamma}^{*})\frac{\pi-R_{i}}{\pi}\right]\right\}+o_{p}(1).$
From Corollary 1.1 and Corollary S1.1, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ if and only if
$2\boldsymbol{\Sigma}_{Y,\widehat{Y}}\succ\frac{1}{1-\pi}\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}$
whereas $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ if and only if
$2\boldsymbol{\Sigma}_{Y,\widehat{Y}}\succ\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}.$
Therefore, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. Moreover, the difference between asymptotic variance-covariance can be derived from Corollary 1.1 and S1.1 as
$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{PPI}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})=\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})-\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}})$
$=\left(\frac{1}{\pi}-1\right){\bf{A}}\left(2\boldsymbol{\Sigma}_{Y,\widehat{Y}}-\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}\right){\bf{A}}-\frac{1}{\pi}{\bf{A}}\left(2\boldsymbol{\Sigma}_{Y,\widehat{Y}}-\frac{1}{1-\pi}\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}\right){\bf{A}}$
$={\bf{A}}\left(\frac{2-\pi}{1-\pi}\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}-2\boldsymbol{\Sigma}_{Y,\widehat{Y}}\right){\bf{A}}.$
Note that if $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is less efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ then $\frac{1}{1-\pi}\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}\succ 2\boldsymbol{\Sigma}_{Y,\widehat{Y}}$, implying
$\frac{2-\pi}{1-\pi}\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}\succ 2\boldsymbol{\Sigma}_{Y,\widehat{Y}}.$
This confirms that when $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is less efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, it will also be less efficient than $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$. A sufficient condition for $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ to be more efficient than $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is
$\mathbb{E}[\epsilon_{Y}^{2}\mid{\bf{X}}]<\mathbb{E}[(\epsilon_{\widehat{Y}}-\epsilon_{Y})^{2}\mid{\bf{X}}]+\frac{1}{1-\pi}\mathbb{E}[\epsilon_{\widehat{Y}}^{2}\mid{\bf{X}}].$
Appendix C Asymptotic properties of $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$
C.1 Proof of Theorem 2
Theorem 2 (Influence function expansion of $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$)
Following Chen and Chen (2000),
$n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{CC}}}-\bm{\beta}^{*}\right)=n^{-\frac{1}{2}}\sum_{i=1}^{n}{\bf{IF}}^{\text{\tiny{CC}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})+o_{p}(1)$
where ${\bf{IF}}^{\text{\tiny{CC}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})={\bf{A}}\left[\bm{\phi}(Y_{i},{\bf{X}}_{i};\bm{\beta}^{*})\frac{R_{i}}{\pi}+\bm{\omega}^{\text{\tiny{CC}}}(\widehat{Y}_{i},{\bf{X}}_{i})\frac{\pi-R_{i}}{\pi}\right]$ and $\bm{\omega}^{\text{\tiny{CC}}}(\widehat{Y},{\bf{X}})$ is defined in (1). Therefore, $n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{CC}}}-\bm{\beta}^{*}\right)$ converges weakly to a zero-mean multivariate normal distribution with variance-covariance
$\hskip 2.84526pt\mathbb{E}\left[{\bf{IF}}^{\text{\tiny{CC}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})\hskip 2.84526pt{\bf{IF}}^{\text{\tiny{CC}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})^{{\sf\scriptscriptstyle{T}}}\right]\hskip 2.84526pt.$
Proof  The proof follows from Chen and Chen (2000), but is detailed for completeness. Recall that
$\begin{split}\widehat{{\bf{W}}}^{\text{\tiny{CC}}}&=\widehat{{\bf{A}}}\left\{\widehat{\text{Cov}}\left[\phi(Y,{\bf{X}};\bm{\beta}^{*}),\phi(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})\right]\right\}\{\widehat{\text{Var}}[\phi(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})]\}^{-1}\widehat{{\bf{A}}}^{-1}\\ &=\widehat{{\bf{A}}}\widehat{\boldsymbol{\Sigma}}_{Y,\widehat{Y}}\widehat{\boldsymbol{\Sigma}}^{-1}_{\widehat{Y},\widehat{Y}}\widehat{{\bf{A}}}^{-1}\end{split}$
where
$\begin{split}&\widehat{{\bf{A}}}=n_{\text{\tiny{lab}}}\left(\sum_{i=1}^{n_{\text{\tiny{lab}}}}{\bf{X}}_{i}{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\right)^{-1},\\ &\widehat{\boldsymbol{\Sigma}}_{Y,\widehat{Y}}=n_{\text{\tiny{lab}}}^{-1}\sum_{i=1}^{n_{\text{\tiny{lab}}}}\left(Y_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\hat{\bm{\beta}}^{\text{\tiny{lab}}}\right)\left(\hat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\hat{\bm{\gamma}}^{\text{\tiny{lab}}}\right){\bf{X}}_{i}{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}},\\ &\mbox{and }\widehat{\boldsymbol{\Sigma}}_{\widehat{Y},\widehat{Y}}=n_{\text{\tiny{lab}}}^{-1}\sum_{i=1}^{n_{\text{\tiny{lab}}}}\left(\hat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\hat{\bm{\gamma}}^{\text{\tiny{lab}}}\right)^{2}{\bf{X}}_{i}{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}.\end{split}$
It follows that
$n^{\frac{1}{2}}\left[\widehat{{\bf{W}}}^{\text{\tiny{CC}}}(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\bm{\gamma}^{*})\right]={\bf{A}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}\boldsymbol{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\left\{n^{-\frac{1}{2}}\sum_{i=1}^{n}\left[{\bf{X}}_{i}\left(\widehat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma}^{*}\right)\right]\frac{R_{i}}{\pi}\right\}+o_{p}(1)$
and
$n^{\frac{1}{2}}\left[\widehat{{\bf{W}}}^{\text{\tiny{CC}}}(\hat{\bm{\gamma}}^{\text{\tiny{all}}}-\bm{\gamma}^{*})\right]={\bf{A}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}\boldsymbol{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\left\{n^{-\frac{1}{2}}\sum_{i=1}^{n}\left[{\bf{X}}_{i}\left(\widehat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma}^{*}\right)\right]\right\}+o_{p}(1).$
Consequently,
$n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{CC}}}-\bm{\beta}^{*}\right)=$ $n^{-\frac{1}{2}}\sum_{i=1}^{n}{\bf{A}}\left[\bm{\phi}(Y_{i},{\bf{X}}_{i};\bm{\beta}^{*})\frac{R_{i}}{\pi}+\bm{\omega}^{\text{\tiny{CC}}}(\widehat{Y}_{i},{\bf{X}}_{i})\frac{\pi-R_{i}}{\pi}\right]+o_{p}(1).$
$=$ $n^{-\frac{1}{2}}\sum_{i=1}^{n}{\bf{IF}}^{\text{\tiny{CC}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})+o_{p}(1).$
The standard CLT implies $n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{CC}}}-\bm{\beta}^{*}\right)$ converges weakly to a zero-mean multivariate normal distribution with variance-covariance
$\mathbb{E}\left[{\bf{IF}}^{\text{\tiny{CC}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})\hskip 2.84526pt{\bf{IF}}^{\text{\tiny{CC}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})^{{\sf\scriptscriptstyle{T}}}\right].$
C.2 Proof of Corollary 2.1
Corollary 2.1 ($\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$)
The difference in the asymptotic covariance of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is
$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{CC}}})=\left(\frac{1}{\pi}-1\right){\bf{A}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}\boldsymbol{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}{\bf{A}}\succeq{\bf{0}}.$
Therefore, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is guaranteed to be at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
Proof  From the proof for Corollary 1.1, $n[\text{aVar}(\hat{\bm{\beta}}^{\text{\tiny{lab}}})]=\frac{1}{\pi}{\bf{A}}\boldsymbol{\Sigma}_{Y,Y}{\bf{A}}.$ By Theorem 2.1, $n[\text{aVar}(\hat{\bm{\beta}}^{\text{\tiny{CC}}})]$ is given by
$\begin{split}&\mathbb{E}\left\{\left[{\bf{A}}\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})\right]^{\otimes 2}\frac{R^{2}}{\pi^{2}}\right\}+\mathbb{E}\left\{\left[{\bf{A}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}\boldsymbol{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})\right]^{\otimes 2}\frac{(\pi-R)^{2}}{\pi^{2}}\right\}\\ +&2\mathbb{E}\left\{\left[{\bf{A}}\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})\right]\left[{\bf{A}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}\boldsymbol{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})\right]^{{\sf\scriptscriptstyle{T}}}\frac{R(\pi-R)}{\pi^{2}}\right\}\\ =&\frac{1}{\pi}{\bf{A}}\boldsymbol{\Sigma}_{Y,Y}{\bf{A}}+\frac{1-\pi}{\pi}{\bf{A}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}\boldsymbol{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}{\bf{A}}-2\frac{1-\pi}{\pi}{\bf{A}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}\boldsymbol{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}{\bf{A}}\\ =&\frac{1}{\pi}{\bf{A}}\boldsymbol{\Sigma}_{Y,Y}{\bf{A}}-\frac{1-\pi}{\pi}{\bf{A}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}\boldsymbol{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}{\bf{A}}.\end{split}$
Hence, $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{CC}}})=\frac{1-\pi}{\pi}{\bf{A}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}\boldsymbol{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}{\bf{A}}\succeq 0$.
C.3 Proof of Corollary 2.2
Corollary 2.2 ($\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is the most efficient estimator in its class)
Let the mean squared error (MSE) of $\hat{\bm{\beta}}$ for estimating $\bm{\beta}^{*}$ be defined as $\mathbb{E}[(\hat{\bm{\beta}}-\bm{\beta}^{*})^{{\sf\scriptscriptstyle{T}}}(\hat{\bm{\beta}}-\bm{\beta}^{*})]$. Analogously to the well-known result for the best linear predictor (e.g., Bickel and Doksum (2015), page 40), $\widehat{{\bf{W}}}^{\text{\tiny{CC}}}$ is the weight, ${\bf{W}}$, that minimizes the asymptotic MSE of estimators of the form $\hat{\bm{\beta}}^{\text{\tiny{lab}}}-{{\bf{W}}}\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}\right)$ for estimating $\bm{\beta}^{*}$.
Proof  Letting $\hat{\bm{\beta}}_{{\bf{W}}}=\hat{\bm{\beta}}^{\text{\tiny{lab}}}-{\bf{W}}(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}})$, the MSE of estimating $\hat{\bm{\beta}}_{{\bf{W}}}$ is
$\mathbb{Q}({\bf{W}})=\mathbb{E}\left\{\left[\hat{\bm{\beta}}^{\text{\tiny{lab}}}-{\bf{W}}(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}})-\bm{\beta}^{*}\right]^{\otimes 2}\right\}.$
Thus, the optimal choice of ${\bf{W}}$ that minimizes the MSE is
${\bf{W}}^{*}=\mathbb{E}[(\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\bm{\beta}^{*})(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}})]\left\{\mathbb{E}\left[(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}})^{\otimes 2}\right]\right\}^{-1}.$
Note that the first term is the covariance between $\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\bm{\beta}^{*}$ and $\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}$, and the second term is the inverse variance-covariance of $\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}$. Since
$n^{\frac{1}{2}}\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}\right)={\bf{A}}\left[n^{-\frac{1}{2}}\sum_{i=1}^{n}{\bf{X}}_{i}(\widehat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma}^{*})\frac{R_{i}-\pi}{\pi}\right]+o_{p}(1),$
we have $\text{Cov}(\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\bm{\beta}^{*},\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}})=\frac{1-\pi}{\pi}{\bf{A}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}{\bf{A}}$ and $\text{Var}(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}})=\frac{1-\pi}{\pi}{\bf{A}}\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}{\bf{A}}$. Thus,
${\bf{W}}^{*}=\frac{1-\pi}{\pi}{\bf{A}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}{\bf{A}}\left[\frac{1-\pi}{\pi}{\bf{A}}\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}{\bf{A}}\right]^{-1}={\bf{A}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}\boldsymbol{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}{\bf{A}}^{-1}={\bf{W}}^{CC}.$