### v5

# Another look at inference after prediction

Abstract
From structural biology to epidemiology, predictions from machine learning (ML) models increasingly complement costly gold-standard data to enable faster, more affordable, and scalable scientific inquiry. In response, prediction-based (PB) inference has emerged to accommodate statistical analysis using a large volume of predictions together with a small amount of gold-standard data. The goals of PB inference are two-fold: (i) to mitigate bias from errors in predictions and (ii) to improve efficiency relative to classical inference using only the gold-standard data. While early PB inference methods focused on bias, their ability to enhance efficiency remains a focus of ongoing research. We revisit a foundational PB inference method and show that a simple modification can be applied to guarantee provable improvements in efficiency. In doing so, we establish new connections between augmented inverse probability weighted estimators (AIPW) and several recently proposed PB inference methods with a similar focus. The utility of our proposal, which leverages prediction-based outcomes to enhance efficiency, is demonstrated through extensive simulation studies and an application to real data from the UK Biobank. Further, we contextualize PB inference by drawing connections to historical literature from economics and statistics, highlighting how classic methods directly inform this contemporary problem.
Correspondence to:
Jessica Gronsbell
Postal address: 700 University Ave, Toronto, ON, Canada, M5G 1Z5
Email: j.gronsbell@utoronto.ca.
Telephone number: 416-978-3452
Keywords: Machine learning, measurement error, missing data, prediction-based inference, semi-parametric inference
1 Introduction
Researchers are increasingly turning to machine learning (ML) models to obtain predictions for outcomes that are prohibitively time-consuming or expensive to measure directly and then conducting statistical analyses which treat the predictions as observed data. 1, 2, 3, 4 This prediction-based (PB) inference approach, introduced by Wang et al,5 is comprised of two steps outlined in Figure 1. First, the outcome of interest is imputed using a prediction from a pre-trained model. Then, the predictions are used to draw inferences about the parameters of interest, such as the association parameters between the outcome and a set of covariates, together with a small amount of gold-standard data.***PB inference has primarily focused on outcomes that are difficult to measure. Recent work has considered predicted covariates (e.g., Kluger et al6). In contrast to classical inference that relies solely on gold-standard data, PB inference aims to leverage the predictions to improve statistical efficiency without introducing bias.
In the last several years, applications of PB inference have arisen in diverse fields, spanning genetics, biology, and medicine. For example, population biobanks, such as the UK Biobank ($n\approx$ 500K individuals), collect high-throughput genomic data on large populations and offer unprecedented opportunities for genetic discovery. However, many phenotypes are only partially observed due to the time and expense of measurement. ML models are frequently used to impute missing outcomes based on fully observed baseline data to enable downstream association studies with phenotypes of interest. 7, 8, 9, 10, 11, 12, 13, 14 Similarly, in structural biology, AlphaFold provides protein conformation predictions without the need for expensive and labor-intensive experiments. 15, 16 In the analysis of electronic health records data, large language models can rapidly ascertain disease status at an accuracy competitive with manual chart review by trained annotators, facilitating clinical and epidemiological studies. 17, 18, 19
Figure 1: Prediction-based (PB) inference. $Y$ is the outcome of interest, which is only partially observed, $\bf{Z}$} is a set of predictors of $Y$, $\widehat{Y}$ is the prediction of $Y$ based on $\bf{Z}$}, and $\bf{X}$} is a set of covariates. The focus of scientific inquiry is on the association between $Y$ and $\bf{X}$}. Classical statistical inference relies solely on the data on $Y$ and $\bf{X}$} while PB inference incorporates $\widehat{Y}$ to make use of all of the available data. Dark red represents labeled observations while light red represents unlabeled observations.
Given the relevance of PB inference, Motwani and Witten 20 recently revisited two seminal methods in the field. In the context of least squares estimation, they showed that the post-prediction inference (PostPi) method proposed by Wang et al5 provides a consistent estimator for the association between an outcome and a set of covariates only in restrictive scenarios (e.g., when the pre-trained ML model perfectly captures the underlying regression function). This is a significant practical limitation. PB inference methods should guarantee valid inference (i.e., hypothesis tests with controlled type I error and confidence intervals with nominal coverage rates) in the common scenario that the prediction model is imperfect. In contrast to PostPi, the prediction powered inference (PPI) method proposed by Angelopoulos et al 21 yields consistent estimation and valid inference regardless of the quality of the ML model. This finding is critical given that the PPI paper, published just over a year ago in the high-impact journal Science, has already been cited nearly 150 times and widely adopted in applications.
In this paper, we further study the PPI estimator by assessing its statistical efficiency. Our contributions are three-fold. First, we make new contributions to the literature identifying inefficiencies of the PPI method. 22, 23, 24 We derive a set of results that characterize when the PPI estimator gains over classical estimation using labeled data only in the context of Z-estimation. Second, we show that a more efficient estimator can be obtained by adding a simple weight to the PPI estimator. This approach was proposed by Chen & Chen (CC) over 25 years ago in the context of double sampling designs 25 and its application within the PB inference setting is new. Third, we draw important connections between augmented inverse probability weighted (AIPW) estimators, our proposal, and several recently proposed methods. In doing so, we identify settings in which the CC estimator is theoretically guaranteed to improve over comparable approaches. Building on these findings, we also revisit literature from statistics and economics to identify work dating back to the 1960s that bears directly on current developments in PB inference. Overall, our contributions put classical statistical methods in conversation with this modern problem to inform future methodological developments and applications.
We begin by formally introducing the PB inference problem, reviewing the contributions of Motwani and Witten20 and summarizing existing work in Section 2. We then study the asymptotic properties of the PPI estimator, showing that it can be outperformed by classical inference methods and introduce the CC estimator in Section 3. Our key finding is that the CC estimator is optimal with respect to statistical efficiency in the class of estimators to which the PPI estimator belongs. In Section 4, we discuss related literature and provide connections between the CC estimator, other PB inference methods, and AIPW estimators. We also develop a set of results that compare the efficiency of the CC estimator with several PB inference methods and identify scenarios where the CC estimator is theoretically optimal. We then evaluate our findings with extensive simulation studies in Section 5. In Section 6, we investigate the association between android fat mass and clinical and demographic covariates in the South Asian population in the UK Biobank to demonstrate the practical performance of these approaches. We close with additional discussion in Section 7. Overall, our theoretical and numerical results show that the CC estimator offers a balance between robustness to ML model specification and statistical efficiency, making it a preferred choice for use in practice.
2 Problem setting & background
2.1 Notation and Assumptions
We let $Y$ denote the outcome of interest and $\bf{Z}$} its predictors. $Y$ is only partially observed and a pre-trained ML model is available to predict it over the full data given $\bf{Z}$}, which is fully observed. The ML model $\hat{f}:\mathcal{Z}\rightarrow\mathcal{Y}$ maps the predictor space $\mathcal{Z}$ to the outcome space $\mathcal{Y}$ and produces predictions $\widehat{Y}=\hat{f}({\bf{Z}})$. We let ${\bf{X}}=(1,X_{1},\dots,X_{p})^{{\sf\scriptscriptstyle{T}}}$ denote fully observed covariates that are potentially associated with $Y$ and that are the focus of scientific inquiry. $\bf{X}$} may also be predictive of $Y$ and these covariates may or may not be included in $\bf{Z}$}. The available data therefore consists of a labeled dataset and an independent unlabeled dataset, respectively denoted as
$\mathscr{L}=\{(Y_{i},{\bf{Z}}_{i}^{{\sf\scriptscriptstyle{T}}},{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}})^{{\sf\scriptscriptstyle{T}}}\mid i=1,\dots,n_{\text{\tiny{lab}}}\}\quad\mbox{and}\quad\mathscr{U}=\{({\bf{Z}}_{i}^{{\sf\scriptscriptstyle{T}}},{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}})^{{\sf\scriptscriptstyle{T}}}\mid i=n_{\text{\tiny{lab}}}+1,\dots,n\},$
where $n=n_{\text{\tiny{lab}}}+n_{\text{\tiny{unlab}}}$, $n_{\text{\tiny{unlab}}}$ is the number of unlabeled observations and $n_{\text{\tiny{lab}}}/n\to\pi$ as $n\to\infty$ for $\pi\in(0,1)$.
We focus on the case in which $\mathscr{L}$ and $\mathscr{U}$ include independent and identically distributed samples from a common population of interest. Additionally, $\hat{f}$ is assumed to be pre-trained with a dataset that is independent of $\mathscr{L}$ and $\mathscr{U}$ and is therefore treated as “fixed” (i.e., the variation in estimating $\hat{f}$ is not of interest in downstream inference). $\hat{f}$ may be a black box function in that its training data and specification are not be available to the users. Approaches that address uncertainty in the estimation of $\hat{f}$ have recently been considered.26
2.2 Parameter of interest
We focus on Z-estimation and aim to make inferences on the population parameter $\bm{\beta}^{*}$ that is the solution to the estimating equation
$\mathbb{E}[\bm{\phi}(Y,{\bf{X}};\bm{\beta})]=\bm{0}.$
$\bm{\phi}(\cdot)$ is a deterministic function that satisfies conditions C1-C4 in the Appendix. The simplest consistent estimator of $\bm{\beta}^{*}$, which we denote by $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, uses only $\mathscr{L}$ for estimation and is the solution to the estimating equation
$\sum_{i=1}^{n_{\text{\tiny{lab}}}}\bm{\phi}(Y_{i},{\bf{X}}_{i};\bm{\beta})=\bm{0}.$
2.3 Motivation for PB inference
Although $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ is consistent and enables valid inference for $\bm{\beta}^{*}$, it has limited precision when gold-standard data is difficult to collect and $n_{\text{\tiny{lab}}}$ is not large. In contrast, the predictions $\widehat{Y}$ can be computed in the full data, leading to a much larger dataset for analysis. However, relying on $\widehat{Y}$ for inference may result in biased estimation depending on the quality of predictions from $\hat{f}$. For example, in “naive” inference, $\widehat{Y}$ is directly substituted for $Y$ in downstream analysis.5 The parameter $\bm{\gamma}^{*}$ targeted by this approach solves
$\mathbb{E}[\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma})]=\bm{0}.$
The corresponding estimator based on $\mathscr{L}$ and $\mathscr{U}$, denoted as $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$, is the solution to
$\sum_{i=1}^{n}\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma})=\bm{0}.$
As the distribution of $(\widehat{Y},{\bf{X}}^{{\sf\scriptscriptstyle{T}}})^{{\sf\scriptscriptstyle{T}}}$ may not coincide with that of $(Y,{\bf{X}}^{{\sf\scriptscriptstyle{T}}})^{{\sf\scriptscriptstyle{T}}}$, $\bm{\gamma}^{*}$ bears no necessary correspondence with $\bm{\beta}^{*}$ except under restrictive scenarios. For example, $\bm{\gamma}^{*}$ coincides with $\bm{\beta}^{*}$ when $\bf{Z}$} contains $\bf{X}$}, the ML model captures the true regression function $\mathbb{E}(Y\mid{\bf{Z}})$, and the focus of inference is on a parameter from a linear regression model. To overcome these limitations, PB inference methods utilize $\mathscr{L}$ and $\mathscr{U}$ to essentially balance the relative contributions of $Y$ and $\widehat{Y}$ in making inference on $\bm{\beta}^{*}$.
2.4 PB inference methods
In the context of least squares estimation, Motwani and Witten20 revisited two seminal strategies for PB inference. One of these is the PostPI method of Wang et al,5 which was the first modern method proposed for PB inference. PostPI consists of two steps to incorporate $\widehat{Y}$ into inference for $\bm{\beta}^{*}$. First, $\mathscr{L}$ is used to model the relationship between $Y$ and $\widehat{Y}$ to address potential inaccuracies in the predictions. Then, this model is used together with $\mathscr{U}$ to make inference on $\bm{\beta}^{*}$ with either analytical or bootstrap methods. This approach was shown to yield a consistent estimator for $\bm{\beta}^{*}$ under the strong and arguably unrealistic assumption that the ML model captures the true regression function. When this assumption is not met, the PostPI estimator may not be consistent. Consequently, corresponding hypothesis tests may not have controlled type I error rates and confidence intervals may not have adequate coverage.
Motwani and Witten20 also studied the PPI method 21 which was proposed a few years after PostPI. Unlike its predecessor, PPI has the crucial property that inference remains valid regardless of the quality of the predictions. That is, $\hat{f}$ need not capture the true regression function for the estimator to maintain consistency for $\bm{\beta}^{*}$. The PPI estimator utilizes a common statistical technique that augments an initial consistent estimator for the target parameter based solely on $\mathscr{L}$ with an additional term that is estimable from both $\mathscr{L}$ and $\mathscr{U}$. The augmentation term converges in probability to zero in order to improve efficiency without introducing bias. This approach dates back to at least Robins et al 27 and has been applied in multiple settings to improve estimation efficiency outside of the PB inference problem. 25, 28, 29, 30, 31 We return to this point in our review of related literature in Section 4.
Formally, the PPI estimator is
$\hat{\bm{\beta}}^{\text{\tiny{PPI}}}=\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}\right)$
where $\hat{\bm{\gamma}}^{\text{\tiny{lab}}}$ solves $\sum_{i=1}^{n_{\text{\tiny{lab}}}}\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma})=\bm{0}$ and $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ solves $\sum_{i=n_{\text{\tiny{lab}}}+1}^{n}\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma})=\bm{0}$Angelopoulos et al21 express $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ as $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}-\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\beta}}^{\text{\tiny{lab}}}\right)$ where $\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\beta}}^{\text{\tiny{lab}}}\right)$ is referred to as a “rectifier.” The rectifier serves the purpose of correcting potential bias in $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ due to inaccuracies in $\widehat{Y}$ for $Y$. We reframe the PPI method to study its statistical efficiency.. In contrast to the PostPI method, Motwani and Witten20 show that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is consistent for $\bm{\beta}^{*}$ and $n^{\frac{1}{2}}(\hat{\bm{\beta}}^{\text{\tiny{PPI}}}-\bm{\beta}^{*})$ converges weakly to a zero-mean multivariate normal distribution regardless of the quality of the prediction. Analogous results for Z-estimation problems are established in the original PPI paper (e.g., see Theorem S1 in the Supplementary Materials). This approach therefore asymptotically controls the type I error rate in hypothesis tests and produces confidence intervals with the nominal coverage level.
However, for a PB inference method to be preferred in practice over classical inference based on $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, it should not only yield valid inference, but also be more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when $\widehat{Y}$ is informative of $Y$ and equally efficient when $\widehat{Y}$ is uninformative of $Y$. Two recent works have empirically identified that the PPI method can be less efficient than classical inference and have proposed extensions of the original method that are theoretically guaranteed to outperform classical inference. The first such extension, called PPI++, abandons the augmentation-based approach of PPI and instead directly augments the estimating equation for $\bm{\beta}^{*}$ with information from $\widehat{Y}$. A tuning parameter is used that controls the amount of information from $\widehat{Y}$ that is used for estimation and is cleverly designed to minimize the trace of the resulting estimator’s asymptotic covariance matrix. Subsequently, Miao et al. 23 proposed Post-Prediction Adaptive inference (PSPA), which adopts a similar strategy, but instead produces an estimator with minimal element-wise variance. While both PPI++ and PSPA offer provable improvements over classical inference using only $\mathscr{L}$, their relationships to the original PPI framework remain underexplored. Furthermore, the conditions under which these methods are optimal have yet to be fully established. We revisit these proposals and clarify some of these connections in Section 4.
3 Efficiency analysis
For the remainder of this paper, we study a modified version of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ given by
$\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}=\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}\right).$
The only difference between $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is that all of the data, rather than the unlabeled data, is used to estimate $\bm{\gamma}^{*}$ in the augmentation term (i.e., $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ is replaced by $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$). When $\pi\approx 0$, $\hat{\bm{\gamma}}^{\text{\tiny{all}}}\approx\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ so that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}\approx\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$. In general, it can be shown that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ gains efficiency over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ whenever $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ gains over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ can be more efficient than $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ when $\widehat{Y}$ is informative of $Y$ (see Appendix, specifically Section B.3, for a detailed analysis). Additionally, estimators of this form are commonly used in the semi-parametric inference literature, which we draw upon in the subsequent section in our asymptotic analysis. We provide general results for Z-estimation and use linear regression as a running example throughout our discussion to further cultivate intuition.
3.1 Asymptotic analysis of the PPI estimator
To study the asymptotic properties of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, we present its influence function expansion in Theorem 1. The detailed derivations are deferred to the Appendix and follow those provided for the original PPI estimator.
Theorem 1 (Influence function expansion of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$).
Let $R_{i}$ be the indicator of whether the $i$th observation is labeled. Then,
$n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}-\bm{\beta}^{*}\right)=n^{-\frac{1}{2}}\sum_{i=1}^{n}{\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})+o_{p}(1),$
where ${\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})={\bf{A}}\left\{\left[\ \bm{\phi}(Y_{i},{\bf{X}}_{i};\bm{\beta}^{*})\right]\frac{R_{i}}{\pi}+\left[\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma}^{*})\right]\frac{\pi-R_{i}}{\pi}\right\}$ and ${\bf{A}}=\mathbb{E}\left[\frac{\partial\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})}{\partial\bm{\beta}}\right]^{-1}$. Therefore, $n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}-\bm{\beta}^{*}\right)$ converges weakly to a zero-mean multivariate normal distribution with variance-covariance $\mathbb{E}\left[{\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})\hskip 2.84526pt{\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})^{{\sf\scriptscriptstyle{T}}}\right].$
We use Theorem 1 to compare the asymptotic variance of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ with that of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. The result in Corollary 1.1 shows that the improvement of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ increases with the proportion of unlabeled data provided that $\widehat{Y}$ is of sufficient quality.
Corollary 1.1 (Comparison of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ with $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$).
Let ${\bf{C}}_{11}=\mathbb{E}\left[\bm{\phi}(Y,X;\bm{\beta}^{*})\bm{\phi}(Y,X;\bm{\beta}^{*})^{{\sf\scriptscriptstyle{T}}}\right]$, ${\bf{C}}_{12}=\mathbb{E}\left[\bm{\phi}(Y,X;\bm{\beta}^{*})\bm{\phi}(\widehat{Y},X;\bm{\beta}^{*})^{{\sf\scriptscriptstyle{T}}}\right]$, ${\bf{C}}_{22}=\mathbb{E}\left[\bm{\phi}(\widehat{Y},X;\bm{\gamma}^{*})\bm{\phi}(\widehat{Y},X;\bm{\gamma}^{*})^{{\sf\scriptscriptstyle{T}}}\right]$, and ${\bf{B}}=\mathbb{E}\left[\frac{\partial\bm{\phi}(\widehat{Y},X;\bm{\gamma}^{*})}{\partial\bm{\gamma}}\right]^{-1}$. The difference in the asymptotic variance-covariance (aVar) of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is
$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})$ $=n\left[\text{aVar}(\hat{\bm{\beta}}^{\text{\tiny{lab}}})-\text{aVar}(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})\right]=\left(\frac{1}{\pi}-1\right)\left(2{\bf{A}}{\bf{C}}_{12}-{\bf{B}}{\bf{C}}_{22}\right){\bf{B}}.$
$\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is therefore more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ if and only if $2{\bf{A}}{\bf{C}}_{12}\succ{\bf{B}}{\bf{C}}_{22}$.
We further break down this result in the case of a linear regression model.
Example 1 (Linear Regression).
Let $\epsilon_{Y}=Y-{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\beta}^{*}$, $\epsilon_{\widehat{Y}}=\widehat{Y}-{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma}^{*}$, $\bm{\Sigma}_{\widehat{Y},\widehat{Y}}=\mathbb{E}\left(\epsilon_{\widehat{Y}}^{2}{\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\right)$, and $\bm{\Sigma}_{Y,\widehat{Y}}=\mathbb{E}\left(\epsilon_{Y}\epsilon_{\widehat{Y}}{\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\right)$. Following Corollary 1.1, the difference in the asymptotic variance-covariance (aVar) of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is
$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})$ $=n\left[\text{aVar}(\hat{\bm{\beta}}^{\text{\tiny{lab}}})-\text{aVar}(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})\right]=\left(\frac{1}{\pi}-1\right){\bf{A}}\left(2\bm{\Sigma}_{Y,\widehat{Y}}-\bm{\Sigma}_{\widehat{Y},\widehat{Y}}\right){\bf{A}}.$
$\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is therefore more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ if and only if $2\bm{\Sigma}_{Y,\widehat{Y}}\succ\bm{\Sigma}_{\widehat{Y},\widehat{Y}}$. Moreover, a sufficient condition for $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ to have lower variance than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ is that
$\mathbb{E}\left[\left(\epsilon_{\widehat{Y}}-\epsilon_{Y}\right)^{2}\mid{\bf{X}}\right]<\mathbb{E}\left(\epsilon_{Y}^{2}\mid{\bf{X}}\right).$
This condition coincides with the intuition that incorporating unlabeled data into estimation can improve efficiency when $\widehat{Y}$ has additional predictive power for $Y$ beyond what is already explained by the regression model of $Y$ on $\bf{X}$}. This condition is similar to the result for mean estimation provided in the Supplementary Materials of Angelopoulos et al 21 and can be used to describe several scenarios when $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is guaranteed to have lower asymptotic variance than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when the ML model captures the true regression function (see Appendix, Section A.4). This insight can be better understood in the case of homoskedasticity.
Specifically, suppose that $\mathbb{E}(\epsilon_{Y}^{2}\mid{\bf{X}})=\sigma_{Y}^{2}$, $\mathbb{E}(\epsilon_{\widehat{Y}}^{2}\mid{\bf{X}})=\sigma_{\widehat{Y}}^{2}$, and $\mathbb{E}(\epsilon_{Y}\epsilon_{\widehat{Y}}\mid{\bf{X}})=\sigma_{Y,\widehat{Y}}$,
$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})=\left(\frac{1}{\pi}-1\right)\left[2\sigma_{Y,\widehat{Y}}-\sigma_{\widehat{Y}}^{2}\right]{\bf{A}}.$
$\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ if and only if
$\rho_{Y,\widehat{Y}}>\frac{1}{2}\frac{\sigma_{\widehat{Y}}}{\sigma_{Y}},$
where $\rho_{Y,\widehat{Y}}=\sigma_{Y,\widehat{Y}}/\sigma_{Y}\sigma_{\widehat{Y}}$ is the partial correlation of the residuals for $Y$ and $\widehat{Y}$ given $\bf{X}$}. Thus, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ gains efficiency over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when $\widehat{Y}$ is a strong predictor for $Y$ after adjusting for $\bf{X}$}, though $\widehat{Y}$ does not necessarily need to be derived from a ML model that captures the true regression function. The improvement is stronger with increasing correlation of the residuals between $Y$ and $\widehat{Y}$ and decreasing variability in the residuals of $\widehat{Y}$.
3.2 Asymptotic analysis of the CC estimator
To further analyze the efficiency of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, we draw on the landmark paper that introduced a class of augmented inverse probability weighted (AIPW) estimators which includes all regular asymptotic linear estimators.27 While connections between the PPI estimator and AIPW have been made in the context of mean estimation21, we explictly leverage this framework to introduce a simple modification of the original PPI estimator with theoretical guarantees of improved efficiency. In the current setting, where the goal is to estimate $\bm{\beta}^{*}$ with $Y$ partially observed and $\widehat{Y}$ fully observed, AIPW estimators have influence functions of the form
${\bf{IF}}(Y,\widehat{Y},{\bf{X}};\bm{\omega})={\bf{A}}\left[\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})\frac{R}{\pi}+\bm{\omega}(\widehat{Y},{\bf{X}})\frac{\pi-R}{\pi}\right],$
with $\bm{\omega}(\widehat{Y},{\bf{X}})$ being a function of $\widehat{Y}$ and $\bf{X}$}. The optimal choice of $\bm{\omega}(\widehat{Y},{\bf{X}})$ with respect to asymptotic variance27 is
$\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})=\mathbb{E}[\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})\mid\widehat{Y},{\bf{X}}].$
By Theorem 1, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is an AIPW estimator and can be seen as an approximation to the optimal estimator with $\bm{\omega}(\widehat{Y},{\bf{X}})=\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})$. This is a good approximation, for example, when $\bf{Z}$} contains $\bf{X}$}, the ML model captures the true regression function, and the focus is in fitting a generalized linear model with the canonical link function, $g(\cdot)$. In this case, $\bm{\phi}(Y,X;\bm{\beta})={\bf{X}}[Y-g(\bm{\beta}^{{\sf\scriptscriptstyle{T}}}{\bf{X}})]$ and $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})=\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})$. However, this is also a scenario where naive inference will yield consistent estimation that is more efficient than inference based on the PPI estimator. More broadly, the PPI estimator is not the most efficient estimator in its class and, from Corollary 1.1, can even be less efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when the quality of the prediction is insufficient. A PB inference strategy should enhance statistical efficiency relative to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ whenever possible and, at the very least, should not lose efficiency from incorporating unlabeled data into estimation.
To improve upon $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, one approach is to assume a parametric model for $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$ that can be estimated from $\mathscr{L}$ for analysis. However, it is often difficult to specify an appropriate model for $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$, particularly when the relationship between $Y$ and $\widehat{Y}$ is unknown. Moreover, when the model for $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$ is misspecified, an efficiency gain is not guaranteed. To overcome these difficulties, Chen and Chen 25 proposed a particular choice of $\bm{\omega}(\widehat{Y},{\bf{X}})$ that is guaranteed to be at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and that does not require modeling of $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$. Their approach is motivated by the double-sampling design which, apart from differences in terminology (i.e., they regard $\widehat{Y}$ as a mismeasured version of $Y$), is analogous to the partially labeled data setting considered here. The CC estimator approximates $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$ with
$\bm{\omega}^{\text{\tiny{CC}}}(\widehat{Y},{\bf{X}})=\text{Cov}[\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*}),\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})]\{\text{Var}[\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})]\}^{-1}\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*}).$ (1)
Relating this estimator back to the PPI estimator, the high-level idea is analogous, but $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$ is approximated by the linear projection of $\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})$ on $\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})$ rather than $\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})$ itself, which effectively protects against potential inadequacies in $\widehat{Y}$ that can lead to efficiency loss.
The corresponding estimator, which we call the CC estimator, can be written as
$\hat{\bm{\beta}}^{\text{\tiny{CC}}}=\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\widehat{{\bf{W}}}^{\text{\tiny{CC}}}\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}\right),$
where $\widehat{{\bf{W}}}^{\text{\tiny{CC}}}$ is a consistent estimator of
${{\bf{W}}}^{\text{\tiny{CC}}}={\bf{A}}{\text{Cov}}[\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*}),\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})]\{{\text{Var}}[\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})]\}^{-1}{\bf{B}}^{-1}$
obtained with $\mathscr{L}$ and $\mathscr{U}$. Therefore, a simple way to ensure that an estimator with the same structure as the PPI estimator has asymptotic variance no greater than that of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ is to add a weight function in front of the augmentation term. We present the influence function expansion for $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ in Theorem 2 to further study its asymptotic variance and compare it with $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$.
Theorem 2 (Influence function expansion of $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$).
Following Chen and Chen 25,
$n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{CC}}}-\bm{\beta}^{*}\right)=n^{-\frac{1}{2}}\sum_{i=1}^{n}{\bf{IF}}^{\text{\tiny{CC}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})+o_{p}(1),$
where ${\bf{IF}}^{\text{\tiny{CC}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})={\bf{A}}\left[\bm{\phi}(Y_{i},{\bf{X}}_{i};\bm{\beta}^{*})\frac{R_{i}}{\pi}+\bm{\omega}^{\text{\tiny{CC}}}(\widehat{Y}_{i},{\bf{X}}_{i})\frac{\pi-R_{i}}{\pi}\right]$ and $\bm{\omega}^{\text{\tiny{CC}}}(\widehat{Y},{\bf{X}})$ is defined in (1). Therefore, $n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{CC}}}-\bm{\beta}^{*}\right)$ converges weakly to a zero-mean multivariate normal distribution with variance-covariance $\hskip 2.84526pt\mathbb{E}\left[{\bf{IF}}^{\text{\tiny{CC}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})\hskip 2.84526pt{\bf{IF}}^{\text{\tiny{CC}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})^{{\sf\scriptscriptstyle{T}}}\right]\hskip 2.84526pt.$
The subsequent corollaries apply Theorem 2 to show our two key results:(1) $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and (2) $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is optimal with respect to efficiency in the class of estimators with a weighted linear augmentation term. These results are formally stated in Corollaries 2.1 and 2.2 and show that the CC method is preferable to the PPI method.
Corollary 2.1 ($\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$).
The difference in the asymptotic variance-covariance of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is
$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{CC}}})=\left(\frac{1}{\pi}-1\right){\bf{A}}{\bf{C}}_{12}{\bf{C}}_{22}^{-1}{\bf{C}}_{12}{\bf{B}}\succeq{\bf{0}}.$
Therefore, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is guaranteed to be at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
Corollary 2.2 ($\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is the most efficient estimator in its class).
Let the mean squared error (MSE) of an estimator $\hat{\bm{\beta}}$ for $\bm{\beta}^{*}$ be defined as $\mathbb{E}[(\hat{\bm{\beta}}-\bm{\beta}^{*})^{{\sf\scriptscriptstyle{T}}}(\hat{\bm{\beta}}-\bm{\beta}^{*})]$. ${\bf{W}}^{\text{\tiny{CC}}}$ is the weight, $\bf{W}$}, that minimizes the MSE of estimators of the form $\hat{\bm{\beta}}^{\text{\tiny{lab}}}-{{\bf{W}}}\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}\right)$ for estimating $\bm{\beta}^{*}$.
We return to the example of linear regression to further break down Corollary 2.1.
Example 1 Continued (Linear Regression). For linear regression, the difference in the asymptotic variance-covariance of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is
$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{CC}}})=\left(\frac{1}{\pi}-1\right){\bf{A}}\bm{\Sigma}_{Y,\widehat{Y}}\bm{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\bm{\Sigma}_{Y,\widehat{Y}}{\bf{A}}\succeq{\bf{0}}$
while
$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})=\left(\frac{1}{\pi}-1\right){\bf{A}}\left(2\bm{\Sigma}_{Y,\widehat{Y}}-\bm{\Sigma}_{\widehat{Y},\widehat{Y}}\right){\bf{A}}.$
In contrast to $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})$, the covariance term, $\bm{\Sigma}_{Y,\widehat{Y}}$, in $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{CC}}})$ is normalized by the residual variance of $\widehat{Y}$, $\bm{\Sigma}_{\widehat{Y},\widehat{Y}}$. This normalization scales the covariance according to the precision of $\widehat{Y}$ and enables $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ to always be at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. Returning to the example of homoskedasticity, we can gain more intuition as to why $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ always improves over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ while $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ does not. In this setting,
$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})$ $=\left(\frac{1}{\pi}-1\right)\left(2\sigma_{Y,\widehat{Y}}-\sigma_{\widehat{Y}}^{2}\right){\bf{A}}$
$=\left(\frac{1}{\pi}-1\right)\left[-\left(\sigma_{Y}\rho_{Y,\widehat{Y}}-\sigma_{\widehat{Y}}\right)^{2}+\sigma_{Y}^{2}\rho_{Y,\widehat{Y}}^{2}\right]{\bf{A}}$
and
$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{CC}}})$ $=\left(\frac{1}{\pi}-1\right)\left(\frac{\sigma_{Y,\widehat{Y}}^{2}}{\sigma_{\widehat{Y},\widehat{Y}}^{2}}\right){\bf{A}}=\left(\frac{1}{\pi}-1\right)\sigma_{Y}^{2}\rho_{Y,\widehat{Y}}^{2}{\bf{A}}.$
$\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is therefore more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ provided that $\rho_{Y,\widehat{Y}}>0$ and the gain in efficiency scales directly with $\rho_{Y,\widehat{Y}}$ and $\sigma_{Y}^{2}$. In contrast, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ will only be more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ provided that $\rho_{Y,\widehat{Y}}>\frac{1}{2}\frac{\sigma_{\widehat{Y}}}{\sigma_{Y}}$ and the gain in efficiency will only scale in this manner with the additional condition that $\rho_{Y,\widehat{Y}}=\frac{\sigma_{\widehat{Y}}}{\sigma_{Y}}$.
4 Related literature
Before evaluating the empirical properties of the CC estimator, we provide a brief overview of several closely related areas. Existing PB inference literature has noted connections with semi-supervised inference, semi-parametric inference, missing data, and measurement error. 21, 22, 23, 32, 26, 33, 34 Here we establish additional connections between augmentation-based estimators and recently proposed PB inference approaches. We also review some modern and classical methods from the economics and statistics literature dating back to the 1960s that directly relate to PB inference. These connections underscore that aspects of this modern problem have been previously examined, albeit with differences in motivation, terminology, and notation. Given the close relationship between the PPI and CC methods, ongoing research in PB inference would benefit from a more comprehensive review of these and other related approaches.
4.1 Extensions of PPI
We first return to the PPI++ and PSPA methods introduced in Section 2.4 that similarly aim to improve the efficiency of the PPI estimator. In contrast to the PPI and CC estimators that augment $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ with information from $\widehat{Y}$, PPI++ and PSPA directly incorporate $\widehat{Y}$ into the estimation equation for $\bm{\beta}^{*}$. Specifically, they aim to find the solution to the augmented estimating equation
$\sum_{i=1}^{n_{\text{\tiny{lab}}}}\bm{\phi}(Y_{i},{\bf{X}}_{i};\bm{\beta})-{\bf{V}}\left\{\sum_{i=1}^{n_{\text{\tiny{lab}}}}\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta})-\sum_{i=1}^{n}\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta})\right\}=0$ (2)
where $\bf{V}$} is a matrix that controls the contribution of $\widehat{Y}$ into estimation of $\bm{\beta}^{*}$Similar to the PPI estimator, the authors propose to use only the unlabeled data in the augmentation term. We provide further discussion of this point in the Appendix.. For PPI++, ${\bf{V}}^{\text{\tiny{PPI++}}}=\lambda{\bf{I}}$, where $\lambda$ is selected to minimize the trace of the variance-covariance matrix of the corresponding estimator. For PSPA, ${\bf{V}}^{\text{\tiny{PSPA}}}=\text{diag}({\bf{v}})$ where $\bf{v}$} is instead selected to minimize element-wise variance. While both of these approaches fall within the AIPW framework by construction, we again return to our example of linear regression to connect these approaches to augmentation-based estimators and, in turn, show that the CC method is preferred over both PSPA and PPI++. We leave further comparisons of the three methods, particularly for general Z-estimation problems, to the Appendix and highlight some of our findings in our numerical studies in Section 5.
Example 1 Continued (Linear Regression). The PSPA and PPI++ estimators are asymptotically equivalent to augmentation-based estimators. That is,
$\hat{\bm{\beta}}^{\text{\tiny{PPI++}}}\approx\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\widehat{{\bf{W}}}^{\text{\tiny{PPI++}}}(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}})\quad\mbox{and}\quad\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}\approx\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\widehat{{\bf{W}}}^{\text{\tiny{PSPA}}}(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}})$
where $\widehat{{\bf{W}}}^{\text{\tiny{PPI++}}}={\bf{A}}\widehat{{\bf{V}}}^{\text{\tiny{PPI++}}}{\bf{A}}^{-1}$, $\widehat{{\bf{W}}}^{\text{\tiny{PSPA}}}={\bf{A}}\widehat{{\bf{V}}}^{\text{\tiny{PSPA}}}{\bf{A}}^{-1}$, and $\widehat{{\bf{V}}}^{\text{\tiny{PPI++}}}$ and $\widehat{{\bf{V}}}^{\text{\tiny{PSPA}}}$ are defined in the Appendix. From Corollary 2.2, the CC estimator is asymptotically at least as efficient as both PPI++ and PSPA. In fact, both can be viewed as special cases of the CC estimator: PSPA restricts $\bf{W}$} to be diagonal, while PPI++ further constrains the diagonal elements to be equal.
4.2 Seemingly unrelated regression
The augmentation-based representation not only unifies the PPI++, PSPA, and CC estimators under a common framework, but also connects to several other estimators developed for linear regression, both historically and in recent work. In particular, the framework underlies seemingly unrelated regression (SUR), a method introduced by economist Arnold Zellner in 1962 for the joint estimation of multiple related regression models. 35, 36, 37. Analogously to the current setting, the regression models have different outcome variables, but may have correlated errors. SUR leverages the correlation between the error terms to produce more efficient estimators compared to estimating each regression separately. More formally, SUR considers the regression models
$Y={\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\beta}_{0}+\epsilon\quad\mbox{and}\quad\widehat{Y}={\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma}_{0}+\epsilon^{\prime}\quad$
where $\mathbb{E}(\epsilon\mid{\bf{X}})=0$ and $\mathbb{E}(\epsilon^{\prime}\mid{\bf{X}})=0.$ An estimator for $\bm{\beta}_{0}$ that is based on augmenting $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ was proposed by Conniffe 38 and is given by:
$\hat{\bm{\beta}}^{\text{\tiny{SUR}}}=\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\widehat{W}^{\text{\tiny{SUR}}}\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}\right)\quad\mbox{where}\quad\widehat{W}^{\text{\tiny{SUR}}}=\frac{\sum_{i=1}^{n_{\text{\tiny{lab}}}}(Y_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\hat{\bm{\beta}}^{\text{\tiny{lab}}})(\widehat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\hat{\bm{\gamma}}^{\text{\tiny{lab}}})}{\sum_{i=1}^{n_{\text{\tiny{lab}}}}(\widehat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\hat{\bm{\gamma}}^{\text{\tiny{lab}}})^{2}}.$
Note that $\widehat{W}^{\text{\tiny{SUR}}}$ is an estimator of $W^{\text{\tiny{SUR}}}=\frac{\mathbb{E}[\text{Cov}(\epsilon,\epsilon^{\prime}\mid{\bf{X}})]}{\mathbb{E}[\text{Var}(\epsilon^{\prime}\mid{\bf{X}})]}$. Similar to the PPI and CC methods, the contribution of the augmentation term to estimation increases with the predictiveness of $\widehat{Y}$ for $Y$. In the extreme case where $\widehat{Y}$ is a perfect predictor for $Y$ such that $\widehat{Y}=Y$, $\widehat{W}^{\text{\tiny{SUR}}}=1$, and $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ is simply $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$. In the homoskedastic case when $\text{Var}(\epsilon\mid{\bf{X}})$ and $\text{Cov}(\epsilon,\epsilon^{\prime}\mid{\bf{X}})$ are constant and do not depend on $\bf{X}$}, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is asymptotically equivalent to $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$. 25 Moreover, when the conditional distribution of $(\epsilon,\epsilon^{\prime})$ given $\bf{X}$} is bivariate normal with mean $\bm{0}$ and has constant variance-covariance, $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ (and thus $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ as well) is the maximum likelihood estimator of $\bm{\beta}_{0}$ and fully efficient under the assumed parametric model. 38 In other scenarios, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ achieves a lower asymptotic variance than $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ due to Corollary 2.2.
The SUR estimator was recently applied in a PB inference problem inspired by genome-wide association studies conducted with population biobanks. These studies often rely on pre-trained models to impute outcomes of interest (i.e., phenotypes or traits) as they are partially missing due to the time and expense of ascertainment. 12 The proposed method, SynSurr, assumes the true and predicted outcomes follow a bivariate normal distribution. The SynSurr estimator is equivalent to $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ and asymptotically equivalent to $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ when $\text{Var}(\epsilon\mid{\bf{X}})$ and $\text{Cov}(\epsilon,\epsilon^{\prime}\mid{\bf{X}})$ are constant. Building on SynSurr, Miao et al 14 proposed the POP-GWAS estimator for a variety of regression models utilizing the augmentation technique. The POP-GWAS estimator is
$\hat{\bm{\beta}}^{\text{\tiny{POP}}}=\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\widehat{W}^{\text{\tiny{POP}}}\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}\right)\quad\mbox{where}\quad\widehat{W}^{\text{\tiny{POP}}}=\frac{n_{\text{\tiny{unlab}}}}{n}\widehat{\text{Corr}}(Y,\widehat{Y}\mid{\bf{X}}).$
In contrast to the SUR estimator, $\hat{\bm{\beta}}^{\text{\tiny{POP}}}$ uses $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ instead of $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$ for the augmentation, and the weight, $\widehat{W}^{\text{\tiny{POP}}}$, is not generally a consistent estimator of $W^{\text{SUR}}$. $\hat{\bm{\beta}}^{\text{\tiny{POP}}}$ is therefore less efficient than $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ when the residuals are bivariate normal with constant variance-covariance and less efficient than $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ in general.
4.3 Augmented estimation for quantifying a treatment effect
The augmented estimation approach has also been widely used for treatment effect estimation.39, 40, 41, 42, 43, 29, 44 Many of these approaches build on methods from survey sampling.45, 46 For example, Tsiatis et al 41 consider covariate adjusted analyses for comparing two treatments in randomized clinical trials. The goal is to estimate the treatment effect $\Delta=\mathbb{E}(Y\mid X=1)-\mathbb{E}(Y\mid X=0)$. A standard estimator of $\Delta$ is $\widehat{\Delta}=\overline{Y}^{1}-\overline{Y}^{0}$, where $\overline{Y}^{1}=n_{1}^{-1}\sum_{i=1}^{n}Y_{i}X_{i}$ and $\overline{Y}^{0}=n_{0}^{-1}\sum_{i=1}^{n}Y_{i}(1-X_{i})$ are the empirical means of $Y$ in the treated and control groups of sizes $n_{1}$ and $n_{0}$, respectively. To utilize information from baseline covariates $\bf{Z}$} for improving efficiency, Tsiatis et al 41 propose an augmented estimator of the treatment effect given by
$\widehat{\Delta}-\sum_{i=1}^{n}(X_{i}-\bar{X})[n_{0}^{-1}h^{(0)}({\bf{Z}}_{i})+n_{1}^{-1}h^{(1)}({\bf{Z}}_{i})],$ (3)
where $\bar{X}=n^{-1}\sum_{i=1}^{n}X_{i}$ and $h^{(k)}({\bf{Z}})$ are arbitrary scalar functions of $\bf{Z}$} for $k=0,1$. As $X$ and $\bf{Z}$} are independent due to randomization, the augmentation term in 3 converges in probability to 0, analogously to the augmentation terms in the PPI and CC estimators. The gain in efficiency similarly arises from the potential correlation between $Y$ and $\bf{Z}$}. This technique has been applied in numerous settings, most recently for the analysis of time-to-event outcomes47 and recurrent events.31
4.4 Control variates in Monte Carlo simulations
Lastly, the augmentation technique has also been utilized to improve the efficiency of Monte Carlo estimation. In 1982, Rothery 48 considered using a control variate, a statistic whose properties are known, to improve precision when estimating power via simulation. Suppose interest lies in estimating the power of a test $T$ of $H_{0}:\theta=0$ for some scalar parameter $\theta$. Let $S$ denote an existing test of the same hypothesis whose power $\mu_{Z}=\mathbb{E}_{\theta}(Z)$ is known. Define $Y$ as an indicator that test $T$ rejects and $Z$ as an indicator that test $S$ rejects. For a set of $n$ simulation replicates, the baseline estimator of power for $T$ is simply $\overline{Y}=n^{-1}\sum_{i=1}^{n}Y_{i}$ while a control variate augmented estimator is given by:
$\frac{1}{n}\sum_{i=1}^{n}\left[Y_{i}-\lambda\{Z_{i}-\mu_{Z}\}\right].$ (4)
The variance of the augmented estimator is minimized at $\lambda^{*}=\frac{\text{Cov}(Y,Z)}{\text{Var}(Z)}$. 48, 49 The augmentation term in 4 reduces the variance of the unadjusted power estimate $\overline{Y}$ in proportion to the correlation between $Y$ and $Z$. Building on work from the survey sampling literature, a similar idea was later applied for semi-supervised estimation of means 50 and for causal effect effect estimation with multiple observational data sources44.
5 Numerical studies
5.1 Data generation
We next evaluate the performance of the CC method and its comparators through extensive simulation studies for both linear and logistic regression. Although the data is generated in a manner similar to prior work in PB inference 22, 51, we consider several additional mechanisms for generating predictions. In all scenarios, we generated four covariates, ${\bf{W}}=(W_{1},W_{2},W_{3},W_{4})$, from a zero-mean multivariate normal distribution with a compound symmetric covariance structure with unit variance and pairwise correlation of 0.4. For linear regression, the true outcome was generated as
$Y=h({\bf{W}})+\epsilon\quad\mbox{where}\quad\epsilon\sim N(0,0.75^{2})$
and $h({\bf{W}})=1+0.1\exp(W_{1})+\sin(W_{2})+W_{3}+W_{3}^{2}+I(W_{4}>0)$. To assess performance of the PB inference methods using different predictions, we generated
$\widehat{Y}=h({\bf{W}})+\tau\quad\mbox{where}\quad\tau\sim N\left(a+b|W_{2}|,\sigma^{2}_{\tau}\right).$
We varied the amount of bias in the prediction by adjusting $a$ and $b$ and the amount of noise by adjusting $\sigma_{\tau}$. The predictions were generated with three types of errors: (i) random error where $a=b=0$ and $\sigma_{\tau}$ is a positive constant, (ii) non-random error where a=2a=-2, $b=0$, and $\sigma_{\tau}$ is a positive constant, and (iii) covariate-dependent error where $a=0$, b=3.5b=-3.5, and $\sigma_{\tau}=c|W_{2}|$ where $c$ is a positive constant. In all three scenarios, $\sigma_{\tau}$ was set to achieve $R^{2}=0.8$ or 0.6 to represent a high and low quality prediction. For logistic regression, we generated the true outcome $Y\sim\text{Bern}\left(\frac{1}{1+\exp[-h({\bf{W}})]}\right).$ The predictions were similarly generated with (i) random error by randomly flipping the true labels with probability $p_{error}$, (ii) non-random error by randomly flipping the labels with $Y=1$ with probability $p_{error}$ and those with $Y=0$ with probability $\frac{p_{error}}{20}$, and (iii) covariate-dependent error by flipping the observations with $Y=1$ with probability $I(W_{4}>0)\times\frac{p_{error}}{4}+I(W_{4}<0)\times 3p_{error}$ and those with $Y=0$ with probability $I(W_{4}>0)\times\frac{p_{error}}{4}+I(W_{4}<0)\times 2p_{error}$. In all three scenarios, $p_{error}$ was set to achieve an overall accuracy of either 0.9 or 0.7 to represent a high and low quality prediction.
For both linear and logistic regression, we focus on making inference for the parameters obtained from regressing $Y$ on ${\bf{X}}=[1,W_{1},W_{3},I(W_{4}>0)]$. The size of the full dataset was $n=10,000$ and the proportion of labeled data was varied by letting $n_{\text{\tiny{lab}}}=300,600,1000$ or $2000$ in each setting. We evaluated the performance of the five aforementioned PB inference methods: $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI++}}}$, and $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$. With the exception of the CC method, all estimators were obtained from the ipd R package.52 In the main text, we present the coverage probability (CP) of the 95% confidence intervals (CIs) to evaluate the quality of inferences and the relative efficiency (RE) to assess the improvement over classical inference for each method. The RE is defined as the ratio of the mean squared error (MSE) of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ relative to the MSE of the given estimator. In the Appendix, we include results for percent bias and comparisons of the asymptotic standard error (ASE) and empirical standard error (ESE). All of our results are summarized over 2,000 replications.
5.2 Results
5.2.1 Linear Regression
For brevity, we present results for the lower quality prediction with $R^{2}=0.6$ and the higher quality prediction with $R^{2}=0.8$ in the Appendix. The general patterns for the PB inference methods are consistent across the settings, but with higher gains in efficiency and more pronounced differences across the methods. Figures 2 - 4 display the results for the three types of errors in the predictions. All methods provide coverage close to the nominal level across all scenarios. With respect to RE, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ generally have significantly higher MSE than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, with $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ offering some improvement over $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ when the proportion of labeled data is larger. In contrast, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI++}}}$, and $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ consistently outperform $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ for all parameters and in all scenarios. For the prediction with random error, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ perform similarly and have REs exceeding 1.5 for all parameters, while $\hat{\bm{\beta}}^{\text{\tiny{PPI++}}}$ is substantially less efficient for the coefficient for $X_{2}$. For the predictions with non-random and covariate-dependent error, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ overall outperforms $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI++}}}$. $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ provides substantial improvement over both methods for the coefficient for $X_{2}$ and over $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ for the coefficients for $X_{1}$ and $X_{3}$. In fact, $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI++}}}$ are both outperformed by $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ for the smaller labeled sample sizes for the coefficient for $X_{2}$.
Figure 2: Coverage probability of the 95% confidence intervals for all methods and relative efficiency for the PB inference methods (PPI, PPIa, PPI++, PSPA, and $C$) for linear regression with a prediction with random error.
Figure 3: Coverage probability of the 95% confidence intervals for all methods and relative efficiency for the PB inference methods (PPI, PPIa, PPI++, PSPA, and $C$) for linear regression with a prediction with non-random error.
Figure 4: Coverage probability of the 95% confidence intervals for all methods and relative efficiency for the PB inference methods (PPI, PPIa, PPI++, PSPA, and $C$) for linear regression with a prediction with covariate-dependent error.
5.2.2 Logistic Regression
We again present results for the lower quality prediction with overall accuracy of 0.7 and the higher quality prediction with 0.9 accuracy in the Appendix. The general patterns are again consistent, but with higher gains in efficiency and more pronounced differences across the methods for the higher quality prediction. Figures 5 - 7 display the results for the three types of errors in the predictions. All methods provide coverage close to the nominal level across all scenarios. With respect to RE, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ underperform relative to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ while $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI++}}}$, and $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ generally have MSE much lower than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ in all scenarios. For the predictions with random error, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI++}}}$, and $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ perform similarly and have REs near 1.2 for all parameters. For the predictions with non-random error, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ outperforms $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI++}}}$ for all parameters, offering an additinal 10-20% improvement in efficiency. For the predictions with covariate-dependent error, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ performs similarly to $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ for all parameters and is substantially more efficient than $\hat{\bm{\beta}}^{\text{\tiny{PPI++}}}$ for the coefficient for $X_{3}$.
Figure 5: Coverage probability of the 95% confidence intervals for all methods and relative efficiency for the PB inference methods (PPI, PPIa, PPI++, PSPA, and $C$) for logistic regression with a prediction with random error.
Figure 6: Coverage probability of the 95% confidence intervals for all methods and relative efficiency for the PB inference methods (PPI, PPIa, PPI++, PSPA, and $C$) for logistic regression with a prediction with non-random error.
Figure 7: Coverage probability of the 95% confidence intervals for all methods and relative efficiency for the PB inference methods (PPI, PPIa, PPI++, PSPA, and $C$) for logistic regression with a prediction with covariate-dependent error.
6 UK Biobank Analysis
We also evaluated the performance of the PB inference methods in an analysis focused on central obesity. Central obesity is excessive accumulation of body fat around the abdominal region and is strongly associated with cardiometabolic diseases such as type 2 diabetes and cardiovascular disease. A major challenge in its study is the limited availability of direct body composition measurements. Android fat mass (AFM), measured via a dual-energy X-ray absorptiometry (DEXA) scan, is a key marker for assessing central obesity in clinical practice. 53 However, due to its high cost, DEXA-derived AFM is usually only measured in a small subset of participants in large-scale studies. For example, in the UK Biobank,54 a prospective cohort study measuring genetic, lifestyle, and health information on approximately 500,000 participants in the UK, DEXA-derived AFM is incompletely ascertained, particularly among populations historically underrepresented in genetic research. 54 For instance, out of $5,397$ South Asian participants in the UK Biobank, only $288$ (5.3%) participants have undergone DEXA scans.
Our goal is to examine associations between AFM and age, sex, and body mass index (BMI) among the South Asian population. Previous research has indicated that this population is at higher risk of central obesity and that it is associated with cardiometabolic diseases at lower BMI thresholds than European populations. 55 Since DEXA-derived AFM is missing for most South Asian participants, it provides a natural opportunity to leverage ML-predicted AFM in downstream analyses. To illustrate the impact of different ML models on the efficiency of the PB inference methods, we trained two random forest models in the White British population ($32,460$ participants) and used them to predict missing AFM values in the South Asian population. We built a “strong” prediction model that incorporated both the covariates of interest (i.e., age, sex, and BMI) as well as additional predictors, including smoking status, alcohol consumption, two bioelectrical impedance measurements of the legs, physical activity scores from the International Physical Activity Questionnaire, and sedentary behavior quantified as total hours spent driving, using a computer, and watching television. We built another “weak” prediction model that included only the additional predictors. The strong prediction model achieved an $R^{2}$ of 0.8 while the weak prediction model had a substantially lower $R^{2}$ of 0.5 in the South Asian subpopulation.
Table 1 presents point estimates, standard errors, and RE, defined as the ratio of the variance of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ to the variance of the given estimator, across the two prediction model settings for five estimators: the baseline estimator ($\hat{\bm{\beta}}^{\text{\tiny{lab}}}$), the PPI estimators ($\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI++}}}$), the PSPA estimator ($\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$), and the CC estimator ($\hat{\bm{\beta}}^{\text{\tiny{CC}}}$). Echoing our theoretical and simulation study results, all methods yield similar point estimates as they are all consistent estimators of the same parameter. In the analysis with the strong prediction model, the predictions provide information on central obesity beyond what is captured by the covariates of interest, satisfying the sufficient condition in Corollary 2.1. Consequently, both $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ outperform $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. In fact, all of the PB inference methods ($\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI++}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$, and $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$) yield substantially smaller standard errors compared to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. Among the PB inference methods, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI++}}}$, and $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ show comparable performance, with average REs of 1.39, 1.39, 1.44, and 1.47, respectively. $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ exhibits higher efficiency gains with an average RE of 1.68. The analysis with the weak prediction model highlights the differences in performance between the PPI and CC methods and reinforces the findings in Corollary 2.1 and Corollary 2.2. Here, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ yield larger standard errors than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ for all parameters of interest, with an average RE of 0.52. In contrast, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$, as well as $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI++}}}$, outperform $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, with average RE values of 1.28, 1.29, and 1.16, respectively. Overall, the CC estimator is the best performing method in terms of efficiency, underscoring its utility in practice.
7 Discussion
In this paper, we expand on existing research into the efficiency of PB inference methods by making the following contributions:
  1. 1.
    We employed the AIPW framework to introduce the CC estimator from the double-sampling literature as a simple modification of the PPI estimator, offering theoretically guaranteed efficiency gains over classical inference. Additionally, we demonstrated that the CC estimator is the most efficient estimator within the class of weighted linear augmentation estimators.
  2. 2.
    We made numerous connections between PB inference and existing literature with direct implications for future methodology and applications. Of note, we provided a set of results relating newly proposed PB inference methods to the CC method. For linear regression, we showed that the PPI++ and PSPA estimators are asymptotically equivalent to weighted linear augmentation estimators and therefore the CC estimator is always at least as efficient as these two methods. This theoretical finding was corroborated in our simulated and real data analyses.
  3. 3.
    We conducted the most extensive simulation study to date comparing existing PB inference approaches. We demonstrated empirically that the CC estimator outperforms the PPI++ and PSPA estimators in numerous simulation scenarios as well as in an analysis of UK Biobank data.
However, our work is not without limitations. First, we focused our analyses on Z-estimation and further work is needed to analyze other statistical procedures. Second, we assumed that the labeled and unlabeled data are drawn from the same underlying distribution or equivalently, that $Y$ is missing completely at random. This assumption may not hold in practice, particularly when samples cannot be randomly selected for labeling due to study design constraints or the use of alternative sampling strategies to enhance efficiency (e.g., stratified sampling). However, the CC estimator can be extended to settings where $Y$ is missing at random by incorporating inverse probability weighting and regression adjustment. Third, our current analysis assumes that only one prediction is available for analysis. In many practical scenarios multiple predictions with unknown accuracy are available to the analyst. The CC estimator can be naturally extended to accommodate multiple predictions and we provide a simulated example in the Appendix. Lastly, we focused on the traditional missing data scenario where the proportion of labeled data, $n_{\text{\tiny{lab}}}/n$, is bounded away from 0. Analyses of semi-supervised settings where $\lim_{n_{\text{\tiny{lab}}}\to\infty}n_{\text{\tiny{lab}}}/n=0$ would be of interest to formally model scenarios where the size of labeled and unlabeled are of different orders of magnitude.
SUPPLEMENTARY MATERIAL
The supplementary materials include proofs for theorems and additional examples for Corollary 1.1.
Code availability:
Our analyses can be replicated with the code at: https://github.com/jlgrons/PBInf-Simulations.
UK Biobank Dataset:
This work used data from the UK Biobank study (https://www.ukbiobank.ac.uk) and our access was approved under application 64875.
ACKNOWLEDGMENTS
J. Gronsbell gratefully acknowledges funding from an NSERC Discovery Grant (RGPIN-2021-03734), the Connaught Fund, the University of Toronto Data Science Institute, and the McLaughlin Center for Genetic Research.