### v6

# Another look at inference after prediction

Abstract
From structural biology to epidemiology, predictions from machine learning (ML) models are increasingly used to complement costly gold-standard data to enable faster, more affordable, and scalable scientific inquiry. In response, prediction-based (PB) inference has emerged to accommodate statistical analysis using a large volume of predictions together with a small amount of gold-standard data. The goals of PB inference are two-fold: (i) to mitigate bias from errors in predictions and (ii) to improve efficiency relative to traditional inference using only the gold-standard data. While early PB inference methods focused on bias, their ability to enhance efficiency remains unclear. We revisit a popular PB inference method and show that a simple modification can be applied to guarantee improvements in efficiency beyond yielding valid inferences when the ML predictions are imperfect. The utility of this approach in leveraging prediction-based outcomes to enhance efficiency is demonstrated through extensive simulation studies and an application to the UK Biobank data. We further contextualize the problem of PB inference through historical literature from economics and statistics to highlight perspectives from classical methods in this contemporary problem.
Correspondence to:
Jessica Gronsbell
Postal address: 700 University Ave, Toronto, ON, Canada, M5G 1Z5
Email: j.gronsbell@utoronto.ca.
Telephone number: 416-978-3452
Keywords: Machine learning, measurement error, missing data, prediction-based inference, semi-parametric inference
1 Introduction
Researchers are increasingly turning to machine learning (ML) models to obtain predictions for outcomes that are prohibitively time-consuming or expensive to measure directly and conducting statistical analyses treating the predictions as observed data. hoffman2024we, fan2024narratives, gao2024semi, miao2024task This prediction-based (PB) inference approach, introduced by Wang et al,wang2020methods is comprised of two steps outlined in Figure 1. First, the outcome of interest is imputed using a prediction from a pre-trained model. Then, the predictions are used to draw inferences about the parameters of interest, such as the association parameters between the outcome and a set of covariates, together with a small amount of gold-standard data.***PB inference has primarily focused on outcomes that are difficult to measure. Recent work has considered predicted covariates (e.g., Kluger et alkluger2025prediction). In contrast to traditional inference that relies solely on gold-standard data, PB inference aims to leverage the predictions to improve statistical efficiency without introducing bias.
In the last several years, applications of PB inference have arisen in diverse fields, spanning genetics, biology, and medicine. For example, population biobanks, such as the UK Biobank ($n\approx$ 500K individuals), collect high-throughput genomic data on large populations and offer unprecedented opportunities for genetic discovery research. However, many traits and disease outcomes are only partially observed due to the time and cost of measurement. ML models are commonly used to impute missing outcomes based on fully observed baseline data to enable downstream association studies. tong2020augmented, alipanahi2021glaucoma, cosentino2023copd, dahl2023phenotype, an2023autocomplete, mccaw2024synthetic, liu2024semiparametric, miao2024popgwas Similarly, in structural biology, AlphaFold provides structural predictions without the need for expensive and labor-intensive experiments for proteomic studies. jumper2021highly, gan2022activation In the analysis of electronic health records data, large language models can rapidly ascertain disease status at an accuracy competitive with manual chart review for epidemiological and clinical studies. yang2022large, zhou2022cancerbert, alsentzer2023zero
Figure 1: Prediction-based (PB) inference. $Y$ is the outcome of interest, which is only partially observed, $\bf{Z}$} is a set of predictors of $Y$, $\widehat{Y}$ is the prediction of $Y$ based on $\bf{Z}$}, and $\bf{X}$} is a set of covariates. The focus of scientific inquiry is on the association between $Y$ and $\bf{X}$}. Traditional statistical inference relies solely on the data on $Y$ and $\bf{X}$} while PB inference incorporates $\widehat{Y}$ to make use of all of the available data. Dark red represents labeled observations while light red represents unlabeled observations.
Given the relevance of PB inference, Motwani and Witten motwani2023revisiting recently revisited two seminal methods in the field. In the context of least squares estimation, they showed that the post-prediction inference (PostPi) method proposed by Wang et alwang2020methods provides a consistent estimator for the association between an outcome and a set of covariates only in restrictive scenarios (e.g., when the pre-trained ML model perfectly captures the underlying regression function). This is a significant practical limitation as it is essential that PB inference methods yield valid inference (i.e., hypothesis tests with controlled type I error and confidence intervals with nominal coverage rates) in the likely scenario that the prediction model is imperfect. In contrast, the prediction powered inference (PPI) method proposed by Angelopoulos et al angelopoulos2023prediction yields consistent estimation and valid inference regardless of the quality of the ML model. This finding is critical given that the PPI paper, published just over a year ago in the high-impact journal Science, has already been cited nearly 150 times and widely adopted in applications.
In this paper, we build on the work of Motwani and Witten motwani2023revisiting and further examine the PPI estimator by assessing its statistical efficiency. Our contributions are two-fold. First, we develop a set of results that characterize when the PPI estimator gains over traditional estimators using labeled data only. We also show that a more efficient estimator can be obtained by adding a simple weight to the PPI estimator. This approach was proposed by Chen & Chen (CC) over 25 years ago in the context of double sampling designs.chen2000unified Second, we revisit literature from statistics and economics and identify work dating back to the 1960s that bears directly on current developments in PB inference. Our review puts classical statistical methods in conversation with this modern problem to inform future methodological advances and applications.
We begin by formally introducing the PB inference problem and reviewing the contributions of Motwani and Wittenmotwani2023revisiting in Section 2. We then study the asymptotic properties of the PPI estimator, showing that it can be outperformed by traditional inference methods, and introduce the CC estimator in Section 3. Our key finding is that the CC estimator is optimal with respect to statistical efficiency in a class of estimators to which the PPI estimator belongs. In Section 4, we discuss related literature and provide connections between the CC estimator and other established methods. We then evaluate our findings with extensive simulation studies in Section 5. In Section 6, we investigate the association between android fat mass and clinical and demographic covariates in the South Asian population in the UK Biobank to demonstrate the practical performance of these approaches. We close with additional discussion in Section 7. Overall, our theoretical and numerical results show that the CC estimator offers a balance between robustness to ML model specification and statistical efficiency, making it a preferred choice for use in practice.
2 Problem setting & background
2.1 Notation and Assumptions
We let $Y$ denote the outcome of interest and $\bf{Z}$} a set of predictors. $Y$ is only partially observed and a pre-trained ML model is available to predict it over the full data given $\bf{Z}$}, which are fully observed. The ML model $\hat{f}:\mathcal{Z}\rightarrow\mathcal{Y}$ maps the predictor space $\mathcal{Z}$ to the outcome space $\mathcal{Y}$ and produces predictions $\widehat{Y}=\hat{f}({\bf{Z}})$. We let ${\bf{X}}=(1,X_{1},\dots,X_{p})^{{\sf\scriptscriptstyle{T}}}$ denote fully observed covariates that are potentially associated with $Y$ and that are the focus of scientific inquiry. $\bf{X}$} may also be predictive of $Y$ and these covariates may or may not be included in $\bf{Z}$}. The available data therefore consists of a labeled dataset and an independent unlabeled dataset, respectively denoted as
$\mathscr{L}=\{(Y_{i},{\bf{Z}}_{i}^{{\sf\scriptscriptstyle{T}}},{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}})^{{\sf\scriptscriptstyle{T}}}\mid i=1,\dots,n_{\text{\tiny{lab}}}\}\quad\mbox{and}\quad\mathscr{U}=\{({\bf{Z}}_{i}^{{\sf\scriptscriptstyle{T}}},{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}})^{{\sf\scriptscriptstyle{T}}}\mid i=n_{\text{\tiny{lab}}}+1,\dots,n\},$
where $n=n_{\text{\tiny{lab}}}+n_{\text{\tiny{unlab}}}$, $n_{\text{\tiny{unlab}}}$ is the number of unlabeled observations, and $n_{\text{\tiny{lab}}}/n\to\pi$ as $n\to\infty$ for $\pi\in(0,1)$.
We focus on the case in which $\mathscr{L}$ and $\mathscr{U}$ include independent and identically distributed samples from a common population of interest. Additionally, $\hat{f}$ is assumed to be pre-trained with a dataset that is independent of $\mathscr{L}$ and $\mathscr{U}$ and is therefore treated as “fixed” (i.e., the variation in estimating $\hat{f}$ is not of interest in downstream inference). $\hat{f}$ may be a black box function in that its training data and specification are not be available to the users. Approaches that address uncertainty in the estimation of $\hat{f}$ have recently been considered.zrnic2024cross
2.2 Parameter of interest
Following Motwani and Witten,motwani2023revisiting we consider least squares estimation and aim to make inferences on the population parameter
$\bm{\beta}^{*}\coloneq\operatorname*{arg\,min}_{\bm{\beta}}\mathbb{E}[(Y-{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\beta})^{2}]=[\mathbb{E}({\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}})]^{-1}\mathbb{E}({\bf{X}}Y),$
which quantifies the association between $Y$ and $\bf{X}$}. The simplest consistent estimator of $\bm{\beta}^{*}$, which we denote by $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, uses only $\mathscr{L}$ for estimation and solves the estimating equation
$\sum_{i=1}^{n_{\text{\tiny{lab}}}}\bm{\phi}(Y_{i},{\bf{X}}_{i};\bm{\beta})\coloneq\sum_{i=1}^{n_{\text{\tiny{lab}}}}{\bf{X}}_{i}(Y_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\bm{\beta})=\bm{0}.$
2.3 Motivation for PB inference
Although $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ is consistent and enables valid inference for $\bm{\beta}^{*}$, it has limited precision when gold-standard data is difficult to collect and $n_{\text{\tiny{lab}}}$ is not large. In contrast, the predictions $\widehat{Y}$ can be computed in the full data, leading to a much larger dataset for analysis. However, relying on $\widehat{Y}$ for inference may result in biased estimation depending on the quality of predictions from $\hat{f}$. For example, in “naive” inference, $\widehat{Y}$ is directly substituted for $Y$ in downstream analysis.wang2020methods The parameter targeted by this approach is
[…]
and the corresponding least squares estimator, denoted as $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$, is the solution to
$\sum_{i=1}^{n}\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma})\coloneq\sum_{i=1}^{n}{\bf{X}}_{i}(\widehat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma})=\bm{0}.$
As the distribution of $(\widehat{Y},{\bf{X}}^{{\sf\scriptscriptstyle{T}}})^{{\sf\scriptscriptstyle{T}}}$ may not coincide with that of $(Y,{\bf{X}}^{{\sf\scriptscriptstyle{T}}})^{{\sf\scriptscriptstyle{T}}}$, $\bm{\gamma}^{*}$ bears no necessary correspondence with $\bm{\beta}^{*}$ except under restrictive scenarios (e.g., when $\bf{Z}$} contains $\bf{X}$} and the ML model captures the true regression function). To overcome these limitations, PB inference methods utilize $\mathscr{L}$ and $\mathscr{U}$ to essentially balance the relative contributions of $Y$ and $\widehat{Y}$ in making inference on $\bm{\beta}^{*}$.
2.4 PB inference methods
Motwani and Wittenmotwani2023revisiting revisited two seminal strategies for PB inference. The first is the PostPI method of Wang et al,wang2020methods which is a two-step procedure to incorporate $\widehat{Y}$ into estimation. First, $\mathscr{L}$ is used to model the relationship between $Y$ and $\widehat{Y}$ to address potential inaccuracies in the predictions. Then, this model is used together with $\mathscr{U}$ to make inference on $\bm{\beta}^{*}$ with either analytical or bootstrap approaches. This approach is shown to yield a consistent estimator for $\bm{\beta}^{*}$ under the strong and unrealistic assumption that the ML model captures the true regression function. When this assumption is not met, the PostPI estimator may not be consistent. Consequently, corresponding hypothesis tests may not have controlled type I error rates and confidence intervals may not have adequate coverage.
Motwani and Wittenmotwani2023revisiting also considered the PPI method,angelopoulos2023prediction which has the crucial property that inference remains valid regardless of the quality of the predictions. That is, the PPI method does not require that $\hat{f}$ capture the true regression function for the estimator to maintain consistency for $\bm{\beta}^{*}$. The proposed estimator utilizes a common statistical technique involving the augmentation of an initial consistent estimator for the parameter of interest with an additional term that converges in probability to 0. This approach dates back to at least Robins et al robins1994estimation and has been applied in multiple settings to improve estimation efficiency outside of the PB inference problem. chen2000unified, tsiatis2006, tian2012covariate, wang2015semiparametric We return to this point in our review of related literature in Section 4.
The PPI estimator is
$\hat{\bm{\beta}}^{\text{\tiny{PPI}}}=\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}\right)$
where $\hat{\bm{\gamma}}^{\text{\tiny{lab}}}$ solves $\sum_{i=1}^{n_{\text{\tiny{lab}}}}\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma})=\bm{0}$ and $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ solves $\sum_{i=n_{\text{\tiny{lab}}}+1}^{n}\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma})=\bm{0}$Angelopoulos et alangelopoulos2023prediction express $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ as $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}-\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\beta}}^{\text{\tiny{lab}}}\right)$ where $\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\beta}}^{\text{\tiny{lab}}}\right)$ is referred to as a “rectifier.” The rectifier serves the purpose of correcting potential bias in $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ due to inaccuracies in $\widehat{Y}$ for $Y$. We reframe the PPI method to study its statistical efficiency.. In contrast to the PostPI method, Motwani and Wittenmotwani2023revisiting show that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is consistent for $\bm{\beta}^{*}$ and $n^{\frac{1}{2}}(\hat{\bm{\beta}}^{\text{\tiny{PPI}}}-\bm{\beta}^{*})$ converges weakly to a zero-mean multivariate normal distribution regardless of the quality of the prediction. This approach therefore asymptotically controls the type I error rate in hypothesis tests and produces confidence intervals with the nominal coverage level.
3 Efficiency analysis
As PB inference methods suppose the availability of at least a small labeled data set, a natural baseline for comparison is the estimator based on $\mathscr{L}$ only, $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. For a PB inference method to be preferred over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, it should not only yield valid inference, but also achieve gains in statistical efficiency relative to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when the prediction is informative of $Y$ and perform no worse when $\widehat{Y}$ is uninformative. For the remainder of this paper, we study the latter property and analyze a modified version of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ given by
$\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}=\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}\right).$
The only difference between $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is that all of the data, rather than the unlabeled data, is used to estimate $\bm{\gamma}^{*}$ in the augmentation term (i.e., $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ is replaced by $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$). When $\pi\approx 0$, $\hat{\bm{\gamma}}^{\text{\tiny{all}}}\approx\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ so that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}\approx\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$. In general, it can be shown that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ gains efficiency over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ whenever $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ gains over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ can be more efficient than $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ when $\widehat{Y}$ is informative of $Y$ (see Appendix, specifically Section B.3, for a detailed analysis). Additionally, estimators of this form are commonly used in the semi-parametric inference literature, which we draw upon in the subsequent section in our asymptotic analysis.
3.1 Asymptotic analysis of the PPI estimator
To study the asymptotic propeties of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, we first derive its influence function expansion in Theorem 1. The detailed derivations are deferred to the Appendix.
Theorem 1 (Influence function expansion of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$).
Let $R_{i}$ be the indicator of whether the $i$th observation is labeled. Then,
$n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}-\bm{\beta}^{*}\right)=n^{-\frac{1}{2}}\sum_{i=1}^{n}{\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})+o_{p}(1),$
where ${\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})={\bf{A}}\left\{\left[\ \bm{\phi}(Y_{i},{\bf{X}}_{i};\bm{\beta}^{*})\right]\frac{R_{i}}{\pi}+\left[\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma}^{*})\right]\frac{\pi-R_{i}}{\pi}\right\}$ and ${\bf{A}}=[\mathbb{E}({\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}})]^{-1}$. Therefore, $n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}-\bm{\beta}^{*}\right)$ converges weakly to a zero-mean multivariate normal distribution with variance-covariance $\mathbb{E}\left[{\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})\hskip 2.84526pt{\bf{IF}}^{\text{\tiny{PPI}}_{a}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})^{{\sf\scriptscriptstyle{T}}}\right].$
We use Theorem 1 to compare the asymptotic variance-covariance of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ with that of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. Corollary 1.1 shows that the benefit of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ scales directly with the predictiveness of $\widehat{Y}$ for $Y$ and inversely with the proportion of labeled data $\pi$.
Corollary 1.1 (Comparison of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ with $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$).
Let $\epsilon_{Y}=Y-{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\beta}^{*}$, $\epsilon_{\widehat{Y}}=\widehat{Y}-{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma}^{*}$, $\bm{\Sigma}_{\widehat{Y},\widehat{Y}}=\mathbb{E}\left(\epsilon_{\widehat{Y}}^{2}{\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\right)$, and $\bm{\Sigma}_{Y,\widehat{Y}}=\mathbb{E}\left(\epsilon_{Y}\epsilon_{\widehat{Y}}{\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\right)$. The difference in the asymptotic variance-covariance (aVar) of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is
$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})$ $=n\left[\text{aVar}(\hat{\bm{\beta}}^{\text{\tiny{lab}}})-\text{aVar}(\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})\right]=\left(\frac{1}{\pi}-1\right){\bf{A}}\left(2\bm{\Sigma}_{Y,\widehat{Y}}-\bm{\Sigma}_{\widehat{Y},\widehat{Y}}\right){\bf{A}}.$
$\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is therefore more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ if and only if $2\bm{\Sigma}_{Y,\widehat{Y}}\succ\bm{\Sigma}_{\widehat{Y},\widehat{Y}}$. A sufficient condition for $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ to have lower variance than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ is that
$\mathbb{E}\left[\left(\epsilon_{\widehat{Y}}-\epsilon_{Y}\right)^{2}\mid{\bf{X}}\right]<\mathbb{E}\left(\epsilon_{Y}^{2}\mid{\bf{X}}\right).$
The sufficient condition in Corollary 1.1 coincides with the intuition that incorporating unlabeled data into estimation can improve efficiency when $\widehat{Y}$ has additional predictive power for $Y$ beyond what is already explained by the regression model of $Y$ on $\bf{X}$}. This condition is similar to the result for mean estimation provided in the Supplementary Materials of Angelopoulos et al angelopoulos2023prediction and can be used to describe several scenarios when $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is guaranteed to have lower asymptotic variance than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when the ML model captures the true regression function (see Appendix, Section A.4). Here we consider a simple example to demonstrate when $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ gains efficiency over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ under homoskedasticity.
Example 1 (Efficiency gain of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ under homoskedasticity).
In the homoskedastic setting where $\mathbb{E}(\epsilon_{Y}^{2}\mid{\bf{X}})=\sigma_{Y}^{2}$, $\mathbb{E}(\epsilon_{\widehat{Y}}^{2}\mid{\bf{X}})=\sigma_{\widehat{Y}}^{2}$, and $\mathbb{E}(\epsilon_{Y}\epsilon_{\widehat{Y}}\mid{\bf{X}})=\sigma_{Y,\widehat{Y}}$,
$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})=\left(\frac{1}{\pi}-1\right)\left[2\sigma_{Y,\widehat{Y}}-\sigma_{\widehat{Y}}^{2}\right]{\bf{A}}.$
$\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ if and only if
$\rho_{Y,\widehat{Y}}>\frac{1}{2}\frac{\sigma_{\widehat{Y}}}{\sigma_{Y}},$
where $\rho_{Y,\widehat{Y}}=\sigma_{Y,\widehat{Y}}/\sigma_{Y}\sigma_{\widehat{Y}}$ is the correlation of the residuals for $Y$ and $\widehat{Y}$ given $\bf{X}$}. Thus $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ gains efficiency over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when $\widehat{Y}$ is a strong predictor for $Y$ after adjusting for $\bf{X}$}, though $\widehat{Y}$ does not necessarily need to be derived from a ML model that captures the true regression function. The improvement is stronger with increasing correlation of the residuals between $Y$ and $\widehat{Y}$ and decreasing variability in the residuals of $\widehat{Y}$.
3.2 Asymptotic analysis of the CC estimator
To further analyze the efficiency of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, we draw on the landmark paper that introduced the class of augmented inverse probability weighted (AIPW) estimators. robins1994estimation In the current setting, where the goal is to estimate $\bm{\beta}^{*}$ with $Y$ partially observed and $\widehat{Y}$ fully observed, these estimators have influence functions of the form
$\text{IF}(Y,\widehat{Y},{\bf{X}};\bm{\omega})={\bf{A}}\left[\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})\frac{R}{\pi}+\bm{\omega}(\widehat{Y},{\bf{X}})\frac{\pi-R}{\pi}\right],$
with $\bm{\omega}(\widehat{Y},{\bf{X}})$ being a function of $\widehat{Y}$ and $\bf{X}$}. The optimal choice of $\bm{\omega}(\widehat{Y},{\bf{X}})$ with respect to asymptotic variance is
$\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})=\mathbb{E}[\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})\mid\widehat{Y},{\bf{X}}].$
By Theorem 1, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is thus an AIPW estimator and can be seen as an approximation to the optimal estimator with $\bm{\omega}(\widehat{Y},{\bf{X}})=\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})$. The PPI estimator is therefore not the most efficient estimator in its class and, from Corollary 1.1, can even be less efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when $\bm{\Sigma}_{\widehat{Y},\widehat{Y}}\succ 2\bm{\Sigma}_{Y,\widehat{Y}}$. A PB inference strategy should enhance statistical efficiency relative to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ whenever possible and, at the very least, should not be negatively impacted by incorporating unlabeled data into estimation. This raises concerns about the practical utility of the PPI estimator.
To improve upon $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, one approach is to assume a parametric model for $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$ that can be estimated from $\mathscr{L}$ for analysis. However, it is often difficult to specify an appropriate model for $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$, particularly when the relationship between $Y$ and $\widehat{Y}$ is unknown. Moreover, when the model for $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$ is misspecified, an efficiency gain is not guaranteed. To overcome these difficulties, Chen and Chen chen2000unified proposed a particular choice of $\bm{\omega}(\widehat{Y},{\bf{X}})$ that is guaranteed to be at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and that does not require modeling of $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$. Their approach is motivated by the double-sampling design which, apart from differences in terminology (i.e., they regard $\widehat{Y}$ as a mismeasured version of $Y$), is analogous to the partially labeled data setting considered here. The CC estimator approximates $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$ with
$\bm{\omega}^{\text{\tiny{CC}}}(\widehat{Y},{\bf{X}})=\text{Cov}[\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*}),\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})]\{\text{Var}[\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})]\}^{-1}\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*}).$ (1)
The corresponding estimator can be written as
$\hat{\bm{\beta}}^{\text{\tiny{CC}}}=\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\widehat{{\bf{W}}}^{\text{\tiny{CC}}}\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}\right),$
where $\widehat{{\bf{W}}}^{\text{\tiny{CC}}}$ is a consistent estimator of
${{\bf{W}}}^{\text{\tiny{CC}}}={\bf{A}}{\text{Cov}}[\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*}),\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})]\{{\text{Var}}[\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})]\}^{-1}{\bf{A}}^{-1}$
obtained with $\mathscr{L}$. Therefore, a simple way to ensure that an estimator with the same structure as the PPI estimator has asymptotic variance no greater than that of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ is to add a weight function in front of the augmentation term. We present the influence function expansion for $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ in Theorem 2 to further study its asymptotic variance and compare it with $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$.
Theorem 2 (Influence function expansion of $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$).
Following Chen and Chen chen2000unified,
$n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{CC}}}-\bm{\beta}^{*}\right)=n^{-\frac{1}{2}}\sum_{i=1}^{n}{\bf{IF}}^{\text{\tiny{CC}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})+o_{p}(1),$
where ${\bf{IF}}^{\text{\tiny{CC}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})={\bf{A}}\left[\bm{\phi}(Y_{i},{\bf{X}}_{i};\bm{\beta}^{*})\frac{R_{i}}{\pi}+\bm{\omega}^{\text{\tiny{CC}}}(\widehat{Y}_{i},{\bf{X}}_{i})\frac{\pi-R_{i}}{\pi}\right]$ and $\bm{\omega}^{\text{\tiny{CC}}}(\widehat{Y},{\bf{X}})$ is defined in (1). Therefore, $n^{\frac{1}{2}}\left(\hat{\bm{\beta}}^{\text{\tiny{CC}}}-\bm{\beta}^{*}\right)$ converges weakly to a zero-mean multivariate normal distribution with variance-covariance $\hskip 2.84526pt\mathbb{E}\left[{\bf{IF}}^{\text{\tiny{CC}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})\hskip 2.84526pt{\bf{IF}}^{\text{\tiny{CC}}}(Y_{i},\widehat{Y}_{i},{\bf{X}}_{i};\bm{\beta}^{*},\bm{\gamma}^{*})^{{\sf\scriptscriptstyle{T}}}\right]\hskip 2.84526pt.$
The subsequent corollaries apply Theorem 2 to show our two key results: (1) $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and (2) $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ has the highest asymptotic efficiency within the class of estimators with a weighted linear augmentation term. These results suggest that the CC method is preferable to the PPI method for gaining efficiency.
Corollary 2.1 ($\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$).
The difference in the asymptotic variance-covariance of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is
$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{CC}}})=\left(\frac{1}{\pi}-1\right){\bf{A}}\bm{\Sigma}_{Y,\widehat{Y}}\bm{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\bm{\Sigma}_{Y,\widehat{Y}}{\bf{A}}\succeq{\bf{0}}.$
Therefore, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is guaranteed to be at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
Corollary 2.2 ($\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is the most efficient estimator in its class).
Let the mean squared error (MSE) of an estimator $\hat{\bm{\beta}}$ for $\bm{\beta}^{*}$ be defined as $\mathbb{E}[(\hat{\bm{\beta}}-\bm{\beta}^{*})^{{\sf\scriptscriptstyle{T}}}(\hat{\bm{\beta}}-\bm{\beta}^{*})]$. Analogously to the well-known result for the best linear predictor (e.g.,Bickel and Doksum,bickel2015mathematical page 40), ${\bf{W}}^{\text{\tiny{CC}}}$ is the weight, $\bf{W}$}, that minimizes the asymptotic MSE of estimators of the form $\hat{\bm{\beta}}^{\text{\tiny{lab}}}-{{\bf{W}}}\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}\right)$ for estimating $\bm{\beta}^{*}$.
From Corollary 2.1, we note that in contrast to $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})$, the covariance term, $\bm{\Sigma}_{Y,\widehat{Y}}$, in $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{CC}}})$ is normalized by the residual variance of $\widehat{Y}$, $\bm{\Sigma}_{\widehat{Y},\widehat{Y}}$. This normalization effectively scales the covariance according to the precision of $\widehat{Y}$ and enables $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ to always be at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. Returning to the previous example, we can gain more intuition as to why $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ possesses this property while $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ does not.
Example 1 continued (Efficiency gain of $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ under homoskedasticity) We revisit the case of homoskedasticity and recall that
$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})$ $=\left(\frac{1}{\pi}-1\right)\left(2\sigma_{Y,\widehat{Y}}-\sigma_{\widehat{Y}}^{2}\right){\bf{A}}$
$=\left(\frac{1}{\pi}-1\right)\left[-\left(\sigma_{Y}\rho_{Y,\widehat{Y}}-\sigma_{\widehat{Y}}\right)^{2}+\sigma_{Y}^{2}\rho_{Y,\widehat{Y}}^{2}\right]{\bf{A}}.$
Theorem 2 can be used to show that
$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{CC}}})$ $=\left(\frac{1}{\pi}-1\right)\left(\frac{\sigma_{Y,\widehat{Y}}^{2}}{\sigma_{\widehat{Y},\widehat{Y}}^{2}}\right){\bf{A}}=\left(\frac{1}{\pi}-1\right)\sigma_{Y}^{2}\rho_{Y,\widehat{Y}}^{2}{\bf{A}}.$
$\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ provided that $\rho_{Y,\widehat{Y}}>0$ and the gain in efficiency scales directly with $\rho_{Y,\widehat{Y}}$ and $\sigma_{Y}^{2}$. In contrast, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ will only be more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ provided that $\rho_{Y,\widehat{Y}}>\frac{1}{2}\frac{\sigma_{\widehat{Y}}}{\sigma_{Y}}$ and the gain in efficiency will only scale in this manner with the additional condition that $\rho_{Y,\widehat{Y}}=\frac{\sigma_{\widehat{Y}}}{\sigma_{Y}}$.
4 Related literature
Before evaluating the empirical properties of the CC estimator, we provide a brief overview of several closely related areas. Existing PB inference literature has established connections with semi-supervised inference, semi-parametric inference, missing data, and measurement error. angelopoulos2023prediction, angelopoulos2023ppi++, miao2023assumption, zrnic2024active, zrnic2024cross, angelopoulos_note_2024, ji2025predictions We identify additional connections between PB inference methods and modern and classical methods from the economics and statistics literature dating back to the 1960s. These connections underscore that aspects of the PB inference problem have been previously examined, albeit with differences in motivation, terminology, and notation. Given the close relationship between the PPI and CC methods, ongoing research in PB inference would benefit from a more comprehensive review of these and other related approaches.
4.1 Seemingly unrelated regression
Seemingly unrelated regression (SUR) was introduced by economist Arnold Zellner in 1962 for simultaneous estimation of multiple related regression models. zellner1962efficient, swamy1975bayesian, schmidt1977estimation The regression models have different outcome variables, but are assumed to have correlated errors. SUR leverages the correlation between the error terms to produce more efficient estimators compared to estimating each regression separately. In the current setting, the SUR problem considers the regression models
$Y={\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\beta}^{*}+\epsilon\quad\mbox{and}\quad\widehat{Y}={\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma}^{*}+\epsilon^{\prime}\quad$
where $\mathbb{E}(\epsilon\mid{\bf{X}})=0$ and $\mathbb{E}(\epsilon^{\prime}\mid{\bf{X}})=0.$ An estimator for $\bm{\beta}^{*}$ that is based on augmenting $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ was proposed by Conniffe conniffe1985estimating and is given by:
$\hat{\bm{\beta}}^{\text{\tiny{SUR}}}=\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\widehat{W}^{\text{\tiny{SUR}}}\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}\right)\quad\mbox{where}\quad\widehat{W}^{\text{\tiny{SUR}}}=\frac{\sum_{i=1}^{n_{\text{\tiny{lab}}}}(Y_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\hat{\bm{\beta}}^{\text{\tiny{lab}}})(\widehat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\hat{\bm{\gamma}}^{\text{\tiny{lab}}})}{\sum_{i=1}^{n_{\text{\tiny{lab}}}}(\widehat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\hat{\bm{\gamma}}^{\text{\tiny{lab}}})^{2}}.$
Note that $\widehat{W}^{\text{\tiny{SUR}}}$ is an estimator of $W^{\text{\tiny{SUR}}}=\frac{\mathbb{E}[\text{Cov}(\epsilon,\epsilon^{\prime}\mid{\bf{X}})]}{\mathbb{E}[\text{Var}(\epsilon^{\prime}\mid{\bf{X}})]}$. Similar to the PPI and CC methods, the contribution of the augmentation term to estimation increases with the predictiveness of $\widehat{Y}$ for $Y$. In the extreme case where $\widehat{Y}$ is a perfect predictor for $Y$ such that $\widehat{Y}=Y$, $\widehat{W}^{\text{\tiny{SUR}}}=1$, and $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ is simply $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$. In the homoskedastic case when $\text{Var}(\epsilon\mid{\bf{X}})$ and $\text{Cov}(\epsilon,\epsilon^{\prime}\mid{\bf{X}})$ are constant and do not depend on $\bf{X}$}, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is asymptotically equivalent to $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$. chen2000unified Moreover, when the conditional distribution of $(\epsilon,\epsilon^{\prime})$ given $\bf{X}$} is bivariate normal with mean $\bm{0}$ and has constant variance-covariance, $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ (and thus $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ as well) is the maximum likelihood estimator of $\bm{\beta}^{*}$ and fully efficient under the assumed parametric model. conniffe1985estimating In other scenarios, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ achieves a lower asymptotic variance than $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$.
The SUR estimator was recently applied in a PB inference problem inspired by genome-wide association studies conducted with population biobanks. These studies often rely on pre-trained models to impute outcomes of interest (i.e., phenotypes or traits) as they are partially missing due to the time and expense of ascertainment. mccaw2024synthetic The proposed method, “SynSurr”, assumes the true and predicted outcomes follow a bivariate normal distribution. The SynSurr estimator is equivalent to $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ and asymptotically equivalent to $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ when $\text{Var}(\epsilon\mid{\bf{X}})$ and $\text{Cov}(\epsilon,\epsilon^{\prime}\mid{\bf{X}})$ are constant. Building on SynSurr, Miao et al miao2024popgwas proposed the “POP-GWAS” estimator for a variety of regression models utilizing the augmentation technique. The POP-GWAS estimator is
$\hat{\bm{\beta}}^{\text{\tiny{POP}}}=\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\widehat{W}^{\text{\tiny{POP}}}\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}\right)\quad\mbox{where}\quad\widehat{W}^{\text{\tiny{POP}}}=\frac{n_{\text{\tiny{unlab}}}}{n}\widehat{\text{Corr}}(Y,\widehat{Y}\mid{\bf{X}}).$
In contrast to the SUR estimator, $\hat{\bm{\beta}}^{\text{\tiny{POP}}}$ uses $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ instead of $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$ for the augmentation, and the weight, $\widehat{W}^{\text{\tiny{POP}}}$, is not generally a consistent estimator of $W^{\text{SUR}}$. $\hat{\bm{\beta}}^{\text{\tiny{POP}}}$ is therefore less efficient than $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ when the residuals are bivariate normal with constant variance-covariance and less efficient than $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ in general.
4.2 Augmented estimation for quantifying a treatment effect
The augmented estimation approach used by the PPI and CC methods has also been widely used for treatment effect estimation.leon2003semiparametric, lu2008improving, tsiatis2008covariate, gilbert2009efficient, zhang2010increasing, tian2012covariate, yang2020combining Many of these approaches build on methods from survey sampling.cassel1976some, sarndal2003model For example, Tsiatis et al tsiatis2008covariate consider covariate adjusted analyses for comparing two treatments in randomized clinical trials. Letting $X$ be an indicator for randomization to the treated or control arms, the focus of inference is on $\beta$ in the model
$\mathbb{E}(Y\mid X)=\alpha+\beta X$
to estimate the treatment effect $\Delta=\mathbb{E}(Y\mid X=1)-\mathbb{E}(Y\mid X=0)$. A standard estimator of $\Delta$ is $\widehat{\Delta}=\overline{Y}^{1}-\overline{Y}^{0}$, where $\overline{Y}^{1}=n_{1}^{-1}\sum_{i=1}^{n}Y_{i}X_{i}$ and $\overline{Y}^{0}=n_{0}^{-1}\sum_{i=1}^{n}Y_{i}(1-X_{i})$ are the empirical means of $Y$ in the treated and control groups of sizes $n_{1}$ and $n_{0}$, respectively. To utilize information from baseline covariates $\bf{Z}$} for improving efficiency, Tsiatis et al tsiatis2008covariate propose an augmented estimator of the treatment effect given by
$\widehat{\Delta}-\sum_{i=1}^{n}(X_{i}-\bar{X})[n_{0}^{-1}h^{(0)}({\bf{Z}}_{i})+n_{1}^{-1}h^{(1)}({\bf{Z}}_{i})],$ (2)
where $\bar{X}=n^{-1}\sum_{i=1}^{n}X_{i}$ and $h^{(k)}({\bf{Z}})$ are arbitrary scalar functions of $\bf{Z}$} for $k=0,1$. As $X$ and $\bf{Z}$} are independent due to randomization, the augmentation term in 2 converges in probability to 0, analogously to the augmentation terms in the PPI and CC estimators. The gain in efficiency similarly arises from the potential correlation between $Y$ and $\bf{Z}$}. This technique has been applied in numerous settings, most recently for the analysis of time-to-event outcomeszhang2025unified and recurrent events.gronsbell2024nonparametric
4.3 Control variates in Monte Carlo simulations
The augmentation technique has also been utilized to improve the efficiency of Monte Carlo estimation. Specifically, Rothery rothery1982 considers using a control variate, a statistic whose properties are known, to improve precision when estimating power via simulation. Suppose interest lies in estimating the power of a test $T$ of $H_{0}:\theta=0$ for some scalar parameter $\theta$. Let $S$ denote an existing test of the same hypothesis whose power $\mu_{Z}=\mathbb{E}_{\theta}(Z)$ is known. Define $Y$ as an indicator that test $T$ rejects and $Z$ as an indicator that test $S$ rejects. For a set of $n$ simulation replicates, the baseline estimator of power for $T$ is simply $\overline{Y}=n^{-1}\sum_{i=1}^{n}Y_{i}$ while a control variate augmented estimator is given by:
$\frac{1}{n}\sum_{i=1}^{n}\left[Y_{i}-\lambda\{Z_{i}-\mu_{Z}\}\right].$ (3)
The variance of the augmented estimator is minimized at $\lambda^{*}=\frac{\text{Cov}(Y,Z)}{\text{Var}(Z)}$. rothery1982, davidson1992 The augmentation term in 3 reduces the variance of the unadjusted power estimate $\overline{Y}$ in proportion to the correlation between $Y$ and $Z$. Building on work from the survey sampling literature, a similar idea was later applied for semi-supervised estimation of means.zhang2019semi
5 Numerical studies
5.1 Data generation
We next consider two primary scenarios to evaluate the performance of the PB inference methods and to demonstrate the implications of our theoretical results. In both settings, the generative model for the outcome was
$Y=\theta_{X_{1}}X_{1}+\theta_{\text{\tiny{INT}}}X_{1}W_{1}+{\bf{W}}^{\top}\bm{\theta}+\epsilon$
where ${\bf{W}}=(W_{1},\dots,W_{q})^{{\sf\scriptscriptstyle{T}}}$ and the error term $\epsilon$ was simulated independently from a standard normal distribution. We generated $({\bf{W}}^{\top},X_{1})^{{\sf\scriptscriptstyle{T}}}$ from a mean-zero multivariate normal distribution with a compound symmetric covariance structure with unit variance and pairwise correlation of 0.4. Our objective was to perform inference on the marginal effect of $X_{1}$ on $Y$ in a linear model. Specifically, we target the parameter $\beta_{1}^{*}$, defined as
$(\beta_{0}^{*},\beta_{1}^{*})=\operatorname*{arg\,min}_{\bm{\beta}}\mathbb{E}\left[(Y-(\beta_{0}+\beta_{1}X_{1}))^{2}\right].$
The predictors to train the ML model, $\bf{Z}$}, were constructed based on $\bf{W}$} and $X_{1}$. The scenarios described below were obtained by varying the values of $\theta_{X_{1}}$, $\theta_{\text{\tiny{INT}}}$, and $\bm{\theta}$, as well as the dimension $q$ of $\bf{W}$} and the composition of the predictors $\bf{Z}$}.
Scenario 1 (Strong and Moderate Prediction Models)
We consider two sub-scenarios to illustrate Corollary 1.1, that is, that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when the sufficient condition holds. In both sub-scenarios, we let $q=10$, $\theta_{X_{1}}=1$, $\theta_{\text{\tiny{INT}}}=0$, and $\bm{\theta}=(1,1,0.5,0.5,0.25,{\bf{0}}_{1\times 5})^{{\sf\scriptscriptstyle{T}}}$. In Scenario 1a, we obtain a strong prediction model by regressing $Y$ on ${\bf{Z}}=(X_{1},W_{1},\dots,W_{10})^{{\sf\scriptscriptstyle{T}}}$ while in Scenario 1b we obtain a weaker prediction model by regressing $Y$ on ${\bf{Z}}=(X_{1},W_{5},X_{1}W_{5})^{{\sf\scriptscriptstyle{T}}}$. The $R^{2}$ of $\widehat{Y}$ for $Y$ is 0.9 and 0.6 in Scenarios 1a and 1b, respectively.
Scenario 2 (Weak Prediction Models)
We consider two sub-scenarios to illustrate the results in Corollary 1.1 and Corollary 2.1, which show that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ can be less efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ with a weak prediction model while $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is at least as efficient $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. In both scenarios, we mimic a common scenario where the prediction model does not contain $X_{1}$. In Scenario 2a, we let $q=9$, $\theta_{X_{1}}=1$, $\theta_{\text{\tiny{INT}}}=0$, and $\bm{\theta}=(0.2,0.1\times{\bf{1}}_{1\times 8})^{{\sf\scriptscriptstyle{T}}}$. We obtain the prediction model by regressing $Y$ on ${\bf{Z}}=(W_{1},\dots,W_{9})^{{\sf\scriptscriptstyle{T}}}$. In Scenario 2b, we let $q=3$, $\theta_{X_{1}}=2$, $\theta_{\text{\tiny{INT}}}=0.25$, and $\bm{\theta}=(0.25,0.25,0.25)^{{\sf\scriptscriptstyle{T}}}$. We obtain the prediction model by regressing $Y$ on ${\bf{Z}}=(W_{1},W_{2},W_{3},W_{2}W_{1},W_{3}W_{1},W_{3}W_{2})^{{\sf\scriptscriptstyle{T}}}$. The $R^{2}$ of $\widehat{Y}$ for $Y$ is 0.5 and 0.4 in Scenarios 2a and 2b, respectively.
In all settings, we generated two datasets from the same underlying distribution: (i) a training dataset to train the ML model and (ii) an independent analysis dataset for evaluating the PB inference methods containing a labeled dataset and an unlabeled dataset. We used 10,000 replications of the analysis dataset for each scenario and a single training set as $\hat{f}$ is assumed to be fixed. The size of the training set was 100,000 and the size of the analysis dataset was $n=10,000$. We varied the proportion of labeled data in the analysis dataset by letting $n_{\text{\tiny{lab}}}=300,600,1000$ or $2000$ in each setting.
We evaluated the performance of the PPI and CC methods: $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, and $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$. We also included three comparison methods: $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ to serve as the baseline, $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$ to highlight the limitations of naive inference, and a PB inference method from a concurrent pre-print, the PoSt-Prediction Adaptive (PSPA) method.miao2023assumption The PSPA method is a generalization of the PPI++ method, an extension of the original PPI estimator that adaptively minimizes the trace of the asymptotic variance-covariance of the PPI estimator. The PSPA estimator, denoted as $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$, is designed to achieve element-wise variance reduction relative to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ by incorporating the information from $\widehat{Y}$ directly into the estimating equation for $\bm{\beta}^{*}$. We present the percent bias for each estimator as well as the coverage probability (CP) of the 95% confidence intervals (CIs). For the four PB inference methods ($\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$, and $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$), we also present the relative efficiency (RE) defined as the ratio of the mean squared error (MSE) of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ relative to the MSE of the PB inference estimator.
5.2 Results
Table 1 presents the percent bias of the various estimators across the different scenarios and labeled dataset sizes. With the exception of $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$, the methods exhibit minimal bias as they are all consistent estimators for the target parameter. The limitations of naive inference are particularly apparent in Scenarios 2a and 2b, which do not include $X_{1}$ in ML model training. $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$ exhibits an average percent bias of -47% in Scenario 2a and -64% in Scenario 2b.
Table 2 summarizes the CPs of the 95% CIs for all methods across the various scenarios and labeled dataset sizes. As expected, $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and the four PB inference estimators ($\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$, and $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$) maintain coverage near the nominal level in all settings. In contrast, the 95% CIs based on $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$ achieve adequate coverage only in Scenario 1a, where a strong prediction model yields negligible bias in estimating $\beta_{1}^{*}$. Coverage drops to 92% in Scenario 1b and falls to 0% in Scenarios 2a and 2b, where bias is substantial. These results underscore the risks of relying on naive inference in practice and highlight the necessity of using reliable PB inference methods.
Figure 2 displays the relative efficiency across the different settings. In Scenario 1, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ perform similarly, with RE significantly above 1, indicating that the PB inference methods have lower variance than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. Comparing Scenario 1a and 1b, we note a significant decrease in the gains in statistical efficiency of the PB inference methods due to the weaker prediction model used in Scenario 1b. However, all methods are at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, as the sufficient condition in Corollary 1.1 is satisfied. Additionally, as the labeled set size increases and hence $n_{\text{\tiny{lab}}}/n$ increases, the efficiency gain decreases as expected. $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is always at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ in Scenarios 1a and 1b, and more efficient in Scenario 1a when the labeled dataset size is larger due to the difference in the augmentation terms of these two estimators.
Scenario 2 highlights the performance differences between $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ and $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$, which underscore our findings in Corollary 1.1 and Corollary 2.1. In both sub-scenarios, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ consistently outperforms both $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ and is substantially more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. In Scenario 2a, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ achieves an average RE of 1.27 across the various sample sizes, indicating a significant gain over $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. In Scenario 2b, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ again demonstrates superior RE compared to all other methods and remains more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ offers a modest efficiency gain in Scenario 2a and comparable performance to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ in Scenario 2b. Both $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ are outperformed by $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ in nearly all scenarios. In Scenario 2a, while $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ remains comparable to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, the RE for $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is as low as 0.88 with $n_{\text{\tiny{lab}}}=2000$. Both $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ demonstrate a substantial loses in efficiency in Scenario 2b, with $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ exhibiting a RE of 0.63 with $n_{\text{\tiny{lab}}}=2000$.
Figure 2: Relative efficiency of the PB inference methods compared to baseline estimation with the labeled data ($\hat{\bm{\beta}}^{\text{\tiny{lab}}}$). A ratio greater than 1 indicates that the PB inference method has a smaller mean squared error (MSE) than that of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
6 UK Biobank Analysis
We also evaluated the performance of the PB inference methods in an analysis focused on central obesity. Central obesity is excessive accumulation of body fat around the abdominal region and is strongly associated with cardiometabolic diseases such as type 2 diabetes and cardiovascular disease. A major challenge in its study is the limited availability of direct body composition measurements. Android fat mass (AFM), measured via a dual-energy X-ray absorptiometry (DEXA) scan, is a key marker for assessing central obesity in clinical practice. zhu_dxa-derived_2022 However, due to its high cost, DEXA-derived AFM is usually only measured in a small subset of participants in large-scale studies. For example, in the UK Biobank,sudlow_uk_2015 a prospective cohort study measuring genetic, lifestyle, and health information on approximately 500,000 participants in the UK, DEXA-derived AFM is incompletely ascertained, particularly among populations historically underrepresented in genetic research. sudlow_uk_2015 For instance, out of $5,397$ South Asian participants in the UK Biobank, only $288$ (5.3%) participants have undergone DEXA scans.
Our goal is to examine associations between AFM and age, sex, and body mass index (BMI) among the South Asian population. Previous research has indicated that this population is at higher risk of central obesity and that it is associated with cardiometabolic diseases at lower BMI thresholds than European populations. sniderman_why_2007 Since DEXA-derived AFM is missing for most South Asian participants, it provides a natural opportunity to leverage ML-predicted AFM in downstream analyses. To illustrate the impact of different ML models on the efficiency of the PB inference methods, we trained two random forest models in the White British population ($32,460$ participants) and used them to predict missing AFM values in the South Asian population. We built a “strong” prediction model that incorporated both the covariates of interest (i.e., age, sex, and BMI) as well as additional predictors, including smoking status, alcohol consumption, two bioelectrical impedance measurements of the legs, physical activity scores from the International Physical Activity Questionnaire, and sedentary behavior quantified as total hours spent driving, using a computer, and watching television. We built another “weak” prediction model that included only the additional predictors. The strong prediction model achieved an $R^{2}$ of 0.8 while the weak prediction model had a substantially lower $R^{2}$ of 0.5 in the South Asian subpopulation.
Table 3 presents point estimates, standard errors, and RE, defined as the ratio of the variance of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ to the variance of the given estimator, across the two prediction model settings for five estimators: the baseline estimator ($\hat{\bm{\beta}}^{\text{\tiny{lab}}}$), the PPI estimators ($\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$), the PSPA estimator ($\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$), and the CC estimator ($\hat{\bm{\beta}}^{\text{\tiny{CC}}}$). We do not include naive inference in our analyses due to the limitations demonstrated in our simulation studies. Echoing our theoretical and simulation study results, all methods yield similar point estimates as they are all consistent estimators of the same parameter. In the analysis with the strong prediction model, the predictions provide information on central obesity beyond what is captured by the covariates of interest, satisfying the sufficient condition in Corollary 2.1. Consequently, both $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ outperform $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. In fact, all four PB inference methods ($\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$) yield substantially smaller standard errors compared to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. Among the four PB inference methods, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, and $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ show comparable performance, with average REs of 1.39, 1.39, and 1.47, respectively. $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ exhibits higher efficiency gains with an average RE of 1.68. The analysis with the weak prediction model highlights the differences in performance between the PPI and CC methods and reinforces the findings in Corollary 2.1 and Corollary 2.2. Here, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ yield larger standard errors than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ for all parameters of interest, with an average RE of 0.52. In contrast, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$, as well as $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$, outperform $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, with average RE values of 1.28 and 1.16, respectively. Overall, the CC estimator is the best performing method in terms of efficiency, underscoring its utility in practice.
7 Discussion
In this paper, we bridge classical statistical methods with a modern problem to equip practitioners and methodologists with a rigorous approach for PB inference. We find that incorporating a simple weight in the PPI estimator previously proposed for studies with double-sampling designschen2000unified yields valid inference that is guaranteed to be at least as efficient as estimation using only labeled data. Our work addresses efficiency shortcomings of the original PPI estimator which can undermine its practical value. Alongside theoretical results supporting these findings, our simulation studies and application to UK Biobank data verify that the PPI method can suffer substantial efficiency loss when the ML model is inaccurate. Additionally, we draw important connections to several established approaches from the economics and statistics literature to provide context for ongoing developments in PB inference.
Our work is not without limitations. First, we focused our analyses on least squares estimation following Motwani and Witten.motwani2023revisiting Further work is needed to analyze other statistical procedures (e.g., Z-estimation problems in general). Second, we assumed that the labeled and unlabeled data are drawn from the same underlying distribution or equivalently, that $Y$ is missing completely at random. This assumption may not hold in practice, particularly when samples cannot be randomly selected for labeling due to study design constraints or the use of alternative sampling strategies to enhance efficiency (e.g., stratified sampling). However, the CC estimator can be extended to settings where $Y$ is missing at random by incorporating inverse probability weighting and regression adjustment. Lastly, we focused on the traditional missing data scenario where the proportion of labeled data, $n_{\text{\tiny{lab}}}/n$, is bounded away from 0. Analyses of semi-supervised settings where $\lim_{n_{\text{\tiny{lab}}}\to\infty}n_{\text{\tiny{lab}}}/n=0$ would be of interest to formally model scenarios where the size of labeled and unlabeled are of different orders of magnitude.
SUPPLEMENTARY MATERIAL
The supplementary materials include proofs for theorems and additional examples for Corollary 1.1.
Code availability:
Our analyses can be replicated with the code at: https://github.com/SelinaS37/PBInference.
UK Biobank Dataset:
This work used data from the UK Biobank study (https://www.ukbiobank.ac.uk) and our access was approved under application 64875.
ACKNOWLEDGMENTS
J. Gronsbell gratefully acknowledges funding from an NSERC Discovery Grant (RGPIN-2021-03734), the Connaught Fund, the University of Toronto Data Science Institute, and the McLaughlin Center for Genetic Research.
Labeled set size
Method 300 600 1000 2000
Scenario 1a
   $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$ -0.110 -0.110 -0.110 -0.110
   $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ -0.006 0.014 0.003 -0.020
   $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ -0.029 -0.017 -0.014 -0.001
   $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ -0.028 -0.015 -0.012 -0.005
   $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ -0.029 -0.015 -0.012 -0.005
   $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ -0.025 -0.014 -0.011 -0.004
Scenario 1b
   $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$ -0.245 -0.245 -0.245 -0.245
   $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ 0.034 -0.004 0.010 0.005
   $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ 0.012 -0.038 -0.015 -0.016
   $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ 0.013 -0.036 -0.013 -0.012
   $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ 0.014 -0.035 -0.012 -0.012
   $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ 0.008 -0.035 -0.012 -0.011
Scenario 2a
   $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$ -47.017 -47.017 -47.017 -47.017
   $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ -0.034 -0.027 -0.011 0.001
   $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ 0.025 0.015 0.024 0.018
   $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ 0.023 0.013 0.021 0.015
   $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ -0.002 -0.004 0.005 0.009
   $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ -0.095 -0.048 -0.019 0.003
Scenario 2b
   $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$ -63.899 -63.899 -63.899 -63.899
   $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ -0.039 -0.013 -0.013 -0.006
   $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ 0.019 0.024 -0.008 0.007
   $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ 0.017 0.022 -0.008 0.004
   $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ -0.036 -0.010 -0.018 -0.006
   $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ -0.081 -0.030 -0.025 -0.009
Table 1: Percent biases of the PPI estimators ($\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$), the CC estimator ($\hat{\bm{\beta}}^{\text{\tiny{CC}}}$), the PSPA estimator ($\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$), the baseline estimator ($\hat{\bm{\beta}}^{\text{\tiny{lab}}}$), and the naive estimator ($\hat{\bm{\gamma}}^{\text{\tiny{all}}}$) for $n=10,000$ and various labeled dataset sizes.
Labeled set size
Method 300 600 1000 2000
Scenario 1a
   $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$ 0.947 0.947 0.947 0.947
   $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ 0.951 0.953 0.949 0.951
   $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ 0.947 0.947 0.952 0.949
   $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ 0.947 0.947 0.951 0.947
   $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ 0.944 0.945 0.951 0.946
   $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ 0.949 0.946 0.951 0.948
Scenario 1b
   $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$ 0.916 0.916 0.916 0.916
   $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ 0.949 0.952 0.950 0.949
   $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ 0.948 0.949 0.947 0.947
   $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ 0.947 0.948 0.947 0.946
   $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ 0.946 0.947 0.947 0.945
   $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ 0.949 0.949 0.948 0.946
Scenario 2a
   $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$ 0.000 0.000 0.000 1 0.000
   $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ 0.950 0.948 0.949 0.949
   $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ 0.946 0.946 0.946 0.948
   $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ 0.948 0.953 0.956 0.963
   $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ 0.945 0.948 0.946 0.949
   $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ 0.943 0.944 0.946 0.946
Scenario 2b
   $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$ 0.000 0.000 0.000 0.000
   $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ 0.928 0.933 0.929 0.935
   $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ 0.944 0.950 0.954 0.949
   $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ 0.949 0.957 0.965 0.973
   $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ 0.939 0.945 0.946 0.948
   $\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$ 0.936 0.942 0.943 0.947
Table 2: Coverage probabilities of the 95% confidence intervals based on the PPI estimators ($\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$), the CC estimator ($\hat{\bm{\beta}}^{\text{\tiny{CC}}}$), the PSPA estimator ($\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$), the baseline estimator ($\hat{\bm{\beta}}^{\text{\tiny{lab}}}$), and the naive estimator ($\hat{\bm{\gamma}}^{\text{\tiny{all}}}$) for $n=10,000$ and various labeled dataset sizes.
Strong Prediction Model
Intercept Age Sex BMI
$\hat{\bm{\beta}}^{\text{\tiny{lab}}}$
EST -4.857 -0.003 0.813 0.173
SE 0.3080 0.0041 0.0702 0.0087
RE 1.00 1.00 1.00 1.00
$\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$
EST -4.528 -0.007 0.901 0.165
SE 0.2602 0.0036 0.0628 0.0070
RE 1.40 1.36 1.25 1.54
$\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$
EST -4.541 -0.007 0.897 0.165
SE 0.2596 0.0035 0.0628 0.0069
RE 1.41 1.36 1.26 1.56
$\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$
EST -4.454 -0.009 0.888 0.167
SE 0.2747 0.0036 0.0592 0.0063
RE 1.22 1.36 1.40 1.87
$\hat{\bm{\beta}}^{\text{\tiny{CC}}}$
EST -4.553 -0.006 0.903 0.163
SE 0.2590 0.0034 0.0570 0.0057
RE 1.41 1.53 1.52 2.27
Weak Prediction Model
Intercept Age Sex BMI
$\hat{\bm{\beta}}^{\text{\tiny{lab}}}$
EST -4.857 -0.003 0.813 0.173
SE 0.3080 0.0041 0.0702 0.0087
RE 1.00 1.00 1.00 1.00
$\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$
EST -4.557 -0.006 0.836 0.166
SE 0.4175 0.0060 0.0982 0.0118
RE 0.54 0.48 0.51 0.54
$\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$
EST -4.572 -0.006 0.835 0.166
SE 0.4170 0.0059 0.0981 0.0117
RE 0.55 0.49 0.51 0.54
$\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$
EST -4.771 -0.001 0.835 0.161
SE 0.2978 0.0044 0.0667 0.0070
RE 1.07 0.90 1.11 1.55
$\hat{\bm{\beta}}^{\text{\tiny{CC}}}$
EST -4.790 -0.004 0.822 0.172
SE 0.2941 0.0038 0.0652 0.0068
RE 1.10 1.21 1.16 1.64
Table 3: UK Biobank analysis of android fat mass based on the baseline estimator ($\hat{\bm{\beta}}^{\text{\tiny{lab}}}$), the PPI estimators ($\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$), the PSPA estimator ($\hat{\bm{\beta}}^{\text{\tiny{PSPA}}}$), and the CC estimator ($\hat{\bm{\beta}}^{\text{\tiny{CC}}}$). The strong prediction model has an $R^{2}$ of 0.8, while the weak prediction model has an $R^{2}$ of 0.5. Estimates (EST), standard errors (SE), and relative efficiency (RE) are reported to evaluate performance. RE is defined as the ratio of the variance of $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ to the variance of the given estimator. An RE higher than 1 indicates that a method yields a smaller variance than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.